### v1

# The second class particle in the half-line open TASEP

Abstract.
We consider the second class particle in half-line open TASEP under two different initial conditions with shock discontinuities. The exact formulas for the distribution of the second class particle can be derived by using the color-position symmetry theorem of colored half-space TASEP. We study the asymptotic distribution of second class particles under the constant scaling and KPZ-type scaling.
1. Introduction
1.1. Overview
The totally asymmetric simple exclusion process (TASEP) serves as a fundamental stochastic transport model, finding applications across diverse realms of out-of-equilibrium physics, including biological microsystems and vehicular traffic. Notably, it establishes intriguing theoretical connections with Kardar-Parisi-Zhang universality [Cor12] and random matrix theory [Fer10]. Due to its simplicity, allowing for analytical treatment to a certain extent, TASEP has been a catalyst for numerous pivotal studies in non-equilibrium statistical mechanics. The model unfolds on a linear chain where nodes can host a single particle, and each particle can exclusively hop to an adjacent empty node, characterizing its ”totally asymmetric” nature by permitting hopping in only one direction. Over time, researchers have explored various model variants, incorporating additional features tailored to different physical contexts.
A first important issue is the introduction of the totally asymmetric exclusion process on the positive integers with a single particle source at the origin (we call it half-line open TASEP). Half-line open TASEP has been of interest due to the rich phase diagram that is expected to hold for limiting distributions based on the boundary strength. The work by [Ligg75] reveals a pivotal insight: the long-term behavior of this process undergoes a phase transition. If the particle production rate at the source and the initial density fall below a critical threshold, the stationary measure is a product measure; otherwise, it exhibits spatial correlation. Building upon the methodology presented in [DEHP93], [Gro04] demonstrated that these correlations can be elucidated through a matrix product representation. This realization has yielded various outcomes within the physics literature, encompassing the derivation of hydrodynamic limits, comprehension of phase diagrams, and the establishment of large deviation principles. A substantial body of literature on this subject exists, exemplified by [DMZ18] and related references. The asymptotic analysis of the half-line open TASEP with a step initial condition has been explored in [BBCS18a] (in the equivalent LPP language, also referencing prior works [BR01a, BR01b, SI04]) and [BBCS18b] (in the equivalent FTASEP language). These studies unveil that starting from an empty configuration, the total number of particles in the system at time $t$ exhibits Gaussian fluctuations on the scale of $t^{1/2}$ when the particle production rate at the source is strictly below 1/2. However, when the production rate surpasses 1/2, Tracy-Widom GSE fluctuations [TW96] manifest on the $t^{1/3}$ scale. In the critical case, where the production rate equals 1/2, Tracy-Widom GOE fluctuations [TW96] also emerge on the $t^{1/3}$ scale. A two-dimensional crossover among the GUE [TW94], GOE, and GSE distributions was elucidated through multipoint distribution analysis. The asymptotic analysis of the half-line open TASEP with a stationary initial condition has been conducted in [BFO20] and [BFO22] (in an equivalent LPP language). These works reveal analogues of Baik-Rains distribution [BR00] and the Airy stat process of Baik-Ferrari-Péché [BFP10] in the half-space geometry. Notably, recent advancements have significantly advanced our understanding of the half-line ASEP, an one-parameter generalization of the half-line TASEP. For further insights, one may refer to [BBCW18, BBC20, IMS22, He23].
Among other interesting issues, the second-class particles has been frequently considered in the literature. The second-class particles play a crucial role, particularly in the presence of shock discontinuities, where their position can effectively indicate the location of the shock (see [Ligg99], Chapter 3). Consequently, several investigations have been dedicated to understanding second-class particles. In the scenario of Bernoulli-Bernoulli initial conditions, it has been demonstrated that the fluctuations of second-class particles exhibit Gaussian behavior on a scale of $t^{1/2}$, and the limiting process corresponds to a Brownian motion (see [Ferr92, FF94]). However, when dealing with shocks having non-random initial conditions, the situation differs due to the emergence of typical Kardar-Parisi-Zhang (KPZ) fluctuations from the dynamics. For initial conditions featuring non-random densities $\lambda<\rho$, [FGN19] has determined that the fluctuations of the second-class particle, starting from the origin, follow the difference of two independent random variables with GOE Tracy-Widom distributions on a $t^{1/3}$ scale. And in [FN23], the limit process is characterized by the difference of two independent $\mathrm{Airy}_{1}$ processes. More classes of initial conditions were explored in [BF22], where identities for the one-point distribution of a second-class particle were derived by employing a symmetry theorem within a multi-colored TASEP framework [BB21]. Additionally, asymptotic fluctuations were analyzed under various scaling conditions. More recent works about the shocks in the interacting systems, see e.g. [FN15, FN17, Nej18, FN20, Nej20] and the references therein.
In this paper, we consider the second class particles in the half-line open TASEP. Following [BF22], we will consider two initial conditions (with one shock or two shocks) for half-line open TASEP which can be explored by employing the color-position symmetry property of colored half-space TASEP. The main results of this paper are that we write down the exact distribution of the second particles in terms of the height function of standard half-line open TASEP with step initial condition, and then the asymptotic results of second class particles can be obtained by using the properties of the latter. We introduce our models and explain the asymptotic results below.
1.2. Multi-color half-line open TASEP with one shock or two shocks
In this paper, we will introduce several multi-color TASEPs on $\mathbb{N}$ with open boundary. All of them are continuous time processes with second class particles which start from distinct initial configurations.
[…]
Figure 1. The half-line open TASEP with one-shock initial condition. Upper: half-line open TASEP $\eta_{t}$ with initial configuration $\eta_{0}$; Bottom: half-line open TASEP $\tilde{\eta}_{t}$ with initial configuration $\tilde{\eta}_{0}$.
The first half-line open TASEP has three colors (we denote the first class particle by 1, second class particle by 2 and hole by +$\infty$+ ∞), which is defined on the state-space parametrized by occupation variables $\{(\eta(x))_{x\in\mathbb{Z}_{>0}}\in\{1,2,+\infty\}^{\mathbb{Z}_{>0}}\}$. The initial configuration $\eta_{0}(x)$ has one second class particle which stands between $M_{1}$ holes (on the interval $[1,M_{1}]$) and $M_{2}$ first class particles (on the interval $[M_{1}+2,M_{1}+M_{2}+1]$), the positions larger than $M_{1}+M_{2}+1$ are all filled with holes. See Figure 1 upper. The state $\eta_{t}(x)$ at given time $t\in[0,\infty)$ and position $x\in\mathbb{Z}_{>0}$ evolves according to the following dynamics: a particle $\eta_{t}(x)$ jumps from site $x$ to $x+1$ at exponental rate 1 if $\eta_{t}(x)<\eta_{t}(x+1)$, a first class particle is created at site 1 at rate $\alpha$ if $\eta_{t}(1)>1$. Otherwise, we keep the state unchanged. See Figure 2.
The second half-line open TASEP (we denote by $\tilde{\eta}_{t}$) is similar with the first one, but with a different initial condition. The second initial configuration(we denote by $\tilde{\eta}_{0}$) can be obtained from the first one by replacing the first class particles on the interval $[M_{1}+2,M_{1}+M_{2}+1]$ with the second class particles. See Figure 1 bottom. Obviously, the initial conditions $\eta_{0}$ and $\tilde{\eta}_{0}$ create one shock.
[…]
Figure 2. The transition rates, where $k\in\{2,+\infty\}$, $i<j$ and $i,j\in\{1,2,+\infty\}$.
We extend the one-shock initial configuration to the two-shock case by the following way. In the two-shock case of the first initial configuration $\eta_{0}$ (we denote by $\xi_{0}$), we put $N$ holes on the interval $[1,N]$, $M$ first class particles on the interval $[N+1,N+M]$, $M$ holes on the interval $[N+M+1,N+2M]$, one second class particle at the position $N+2M+1$, $N$ first class particles on the interval $[N+2M+2,2N+2M+1]$, and the positions larger that $22N+2M+$ 2 are all filled with holes. We denote the half-line open TASEP with initial condition $\xi_{0}$ by $\xi_{t}$. See Figure 3 Upper. The dynamics of $\xi_{t}$ are also given by Figure 2.
The two-shock case of the second initial configuration $\tilde{\eta}_{0}$ (we denote by $\tilde{\xi}_{0}$) can be obtained from $\xi_{0}$ by replacing the particles on the interval $[N+1,N+M]$ with the second class particles, the particles on the interval $[N+2M+1,2N+2M+1]$ with the third class particles (we denote by 3). We denote the half-line open TASEP with initial condition $\tilde{\xi}_{0}$ by $\tilde{\xi}_{t}$. See Figure 3 Bottom. The dynamics of $\tilde{\xi}_{t}$ are also given by Figure 2, but with $k\in\{2,3,+\infty\}$, $i<j$ and $i,j\in\{1,2,3,+\infty\}$.
[…]
Figure 3. The half-line open TASEP with two-shock initial condition. Upper: half-line open TASEP $\xi_{t}$ with initial configuration $\xi_{0}$; Bottom: half-line open TASEP $\tilde{\xi}_{t}$ with initial configuration $\tilde{\xi}_{0}$.
1.3. Asymptotic distribution of the second class particles under the constant scaling
In this section, we consider the asymptotic probability distribution of the second class particles by taking the constant $M_{1},M_{2}$. Before stating our main results, let us introduce some notation. For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we denote $\ell(\eta)=\#\{i:\eta_{i}=1\}$ and $A(\eta)=\#\{\text{clusters of 1's in}\ \eta\}$. For $1\leq j\leq A(\eta)$, we denote the number of 1 in the $j$-th cluster of 1’s by $\sigma_{j}(\eta)$. It is easy to see that $\sum_{j=1}^{A(\eta)}\sigma_{j}(\eta)=\ell(\eta)$. For $11\leq j\leq A(\eta)-$ 1, we denote the number of 0 in the cluster of 0’s between the $j$-th and $(j+1)$-th cluster of 1’s by $\tau_{j}(\eta)$. And $\tau_{0}(\eta)$ is the number of 0 before the first cluster of 1’s, $\tau_{A(\eta)}(\eta)$ is the number of 0 after the last cluster of 1’s. Then we have $\sum_{k=0}^{A(\eta)}\tau_{k}(\eta)=L-\ell(\eta)$. We will use the notations $\Psi_{i}(\eta)=\sum_{j=i}^{A(\eta)}\sigma_{j}(\eta)$ for $11\leq i\leq A(\eta)+$ 1, and $\Phi_{j}(\eta)=\sum_{k=j}^{A(\eta)}\tau_{k}(\eta)$ for $10\leq j\leq A(\eta)+$ 1, where $\Psi_{A(\eta)+1}(\eta)=\Phi_{A(\eta)+1}(\eta)=0$. For example, if we take $L=12$ and $\eta=(0,0,1,0,0,1,1,0,1,1,0,0)$, then we have $\ell(\eta)=5$, $A(\eta)=3$ and
$\sigma_{1}(\eta)=1,\sigma_{2}(\eta)=2,\sigma_{3}(\eta)=2;$
$\tau_{0}(\eta)=2,\tau_{1}(\eta)=2,\tau_{2}(\eta)=1,\tau_{3}(\eta)=2;$
$\Psi_{1}(\eta)=5,\Psi_{2}(\eta)=4,\Psi_{3}(\eta)=2,\Psi_{4}(\eta)=0;$
$\Phi_{0}(\eta)=7,\Phi_{1}(\eta)=5,\Phi_{2}(\eta)=3,\Phi_{3}(\eta)=2,\Phi_{4}(\eta)=0.$
Definition 1.1.
For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we define the DEHP-tree of $\eta$ by the following way:
  • start from a node represented by a tuple $(\Psi_{1}(\eta),\Phi_{0}(\eta))$, which we call the root.
[…]
  • When $(x,y)\in\{(x,y):x=\Psi_{i}(\eta)\ \text{and}\ \Phi_{i}(\eta)<y\leq\Phi_{i-1}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)+1\}$, the node $(x,y)$ has only one child. When $(x,y)\in\{(x,y):x=\Psi_{i}(\eta)\ \text{and}\ 1\leq y\leq\Phi_{i}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)\}\cup\{(x,y):\Psi_{i+1}(\eta)<x<\Psi_{i}(\eta),1\leq y\leq\Phi_{i}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)\}$, the node $(x,y)$ has two children.
  • end with the nodes $(k,0)$ where $k=0,1,\cdots,\Psi_{1}(\eta)$ which we call the endpoints.
Note that there exists at least one directed path from the root $(\Psi_{1}(\eta),\Phi_{0}(\eta))$ to any child $(x,y)$, we define the weight of each directed path as the product of weights on that path. Define the partition function of a node $(x,y)$ as the sum of weights of all possible directed paths from the root $(\Psi_{1}(\eta),\Phi_{0}(\eta))$ to that node, we denote it by $Z_{x,y}(\eta)$. In particular, we denote the partition function of endpoints $(k,0)$ as $Z_{k}(\eta)$, where $k=0,1,\cdots,\Psi_{1}(\eta)$. See Figure 4 for illustration.
$(5,7)$$c/\alpha$$(5,6)$$c/\alpha$$(5,5)$$c$$(5,4)$$c$$(4,5)$$c$$(5,3)$$c$$(4,4)$$c$$(5,2)$$c$$(4,3)$$c$$(5,1)$$c$$(4,2)$$c$$c$$(3,3)$$c$$c$$c$$c$$c$$c$$(5,0)$$(4,1)$$(3,2)$$(2,3)$$c$$c$$c$$c$$(4,0)$$(3,1)$$(2,2)$$c$$c$$c$$c$$(3,0)$$(2,1)$$(1,2)$$c$$c$$c$$c$$(2,0)$$(1,1)$$(0,2)$$c$$c$$(1,0)$$(0,1)$$c/\alpha$$c/\alpha$$c/\alpha$$c/\alpha$$c/\alpha$$(0,0)$
Figure 4. The DEHP-tree of $\eta=(0,0,1,0,0,1,1,0,1,1,0,0)$; the DEHP-tree of $\eta$ start from the node $(5,7)$ and end with the nodes $(k,0)$ for $0\leq k\leq 5$; the nodes $(5,7)$, $(5,6)$, $(4,5)$, $(4,4)$, $(2,3)$, $(0,2)$, $(0,1)$ have only one child, other nodes have two children.
Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
Theorem 1.2 (Theorem 4.9).
The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
(1.1) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n},$
when $\alpha\leq\frac{1}{2}$, where we use the binomial coefficient $\binom{p}{k}=\frac{p!}{k!(p-k)!}$; and it is given by
(1.2) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
Let $0\leq m\leq M_{2}$, the asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
(1.3) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(1.4) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ have the following relationships:
(1.5) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=M_{2}+1).$
The asymptotic distribution of the unique second class particle disappearing from $\eta_{t}$ is given by the following formula:
(1.6) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\sum_{m=0}^{M_{2}}\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m).$
Our second main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$:
Theorem 1.3 (Theorem 4.10).
The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
(1.7) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(1.8) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
Let $0\leq s\leq N$, the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
(1.9) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(1.10) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ have the following relationships:
(1.11) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=N+1).$
The asymptotic distribution of the unique second class particle disappearing from $\xi_{t}$ is given by the following formula:
(1.12) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\sum_{s=0}^{N}\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s).$
Remark 1.4.
The asymptotic distribution of the height function of second class particles in $\tilde{\xi}_{t}$ is similar with (1.3) and (1.4). For the sake of brevity, we only provide the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$. The third class particles can be regarded as the second-type second class particles.
1.4. Asymptotic distribution of the second class particles under the KPZ scaling
In this section, we consider the asymptotic probability distribution of the second class particles under the KPZ scaling $M_{1}=M_{2}=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$. Below we will use the half-space ${Airy}_{2}$ process $\mathcal{A}_{\varpi}^{HS}$ which was introduced in [BBCS18a], see section 5 for the definition. Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
Theorem 1.5 (Theorem 5.4).
Assume that $M_{1}=M_{2}=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$. Setting $\alpha=\frac{1+2^{4/3}\varpi t^{-1/3}}{2}$, where $\varpi\in\mathbb{R}$. The asymptotic distribution of the unique second class particle disappearing from the half-space TASEP $\eta_{t}$ is given by:
(1.13) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\mathbb{P}\left(\mathcal{A}_{\varpi}^{HS}(0)+2^{4/3}a^{2}\leq\mathcal{A}_{\varpi}^{HS}(2^{2/3}a)\right).$
The asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
(1.14) $\lim_{t\to\infty}\frac{\mathbf{N}_{2}(1,t)-at^{2/3}}{t^{1/3}}\stackrel{{\scriptstyle d}}{{=}}\min\left\{2^{-4/3}\mathcal{A}_{\varpi}^{HS}(0)+a^{2}-2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{2/3}a),0\right\}.$
Our second main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$:
Theorem 1.6 (Theorem 5.5).
Assume that $M=N=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$. Setting $\alpha=\frac{1+2^{4/3}\varpi t^{-1/3}}{2}$, where $\varpi\in\mathbb{R}$. The asymptotic distribution of the unique second class particle disappearing from the half-space TASEP $\xi_{t}$ is given by:
(1.15) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\mathbb{P}\left(2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{5}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-3a^{2}\right.\\ \left.+\min\left\{2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2},0\right\}\geq 0\right).$
The asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
(1.16) $\lim_{t\to\infty}\frac{\mathsf{N}_{3}(1,t)-at^{2/3}}{t^{1/3}}\stackrel{{\scriptstyle d}}{{=}}\min\left\{0,2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{2/3}a)+3a^{2}-2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{5/3}a)\right.\\ \left.+\max\{0,2^{-4/3}\mathcal{A}_{\varpi}^{HS}(0)+a^{2}-2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{2/3}a)\}\right\}.$
1.5. Outline
The paper is organized as follows. In Section 2 we introduce the colored half-space TASEP and the color-position symmetry theorem. By using the color-position symmetry theorem, we derive the exact distribution of the second class particles in different kinds of multi-color half-line TASEPs in Section 3. In Section 4 and Section 5, we write down the asymptotic results for the second class particles under the constant-scaling and KPZ-scaling.
Acknowledgments
I am grateful to A. Bufetov for very useful discussions. The work was partially supported by A. Bufetov’s Grant from the European Research Council (ERC), Grant Agreement No. 101041499.
2. Color-position symmetry theorem of colored half-space TASEP
In this section, we first define the colored half-space TASEP. Subsequently, we describe the evolution of colored half-space TASEP as the random walk on the type BC Hecke algebra, and introduce the algebraic properties of the colored half-space TASEP. Finally, we will use the algebraic structure of colored half-space TASEP to derive the color-position symmetry theorem.
2.1. Colored half-space TASEP
The colored half-space totally asymmetric simple exclusion process (colored half-space TASEP) is described as a continuous time Markov process operating on the set of bijections from $\mathbb{N}\cup-\mathbb{N}$ to itself. Initially, particles of various colors are placed at corresponding positions, with each particle at position $x$ assigned the color $x$. Formally, the configurations of the system are represented by functions $\eta$, where $\eta(x)$ denotes the color of the particle at position $x$. The process commences with the identity function, $\eta(x)=x$.
The dynamics of this model involve associating exponential clocks of rate 1 with each edge $(x,x+1)$ for $x\geq 1$, and an exponential clock of rate $\alpha$ with the edge $(-1,1)$. When a clock attached to the edge $(x,x+1)$ with rate 1 rings, particles are swapped at positions $x$ and $x+1$ if $\eta(x)<\eta(x+1)$, and a similar swap occurs for particles at positions $-x$ and $-x-1$ at the same time. When the clock with rate $\alpha$ rings, a swap occurs between $\eta(-1)$ and $\eta(1)$ if $\eta(-1)<\eta(1)$. Otherwise, we remain the configuration of system unchanged. See Figure 5 for an illustration. For any fixed time $t$, there is almost surely in finding a sequence $i_{k}\in\mathbb{N}$ such that no clock associated with the edge $(i_{k},i_{k}+1)$ rings up to time $t$. Given this sequence, the process can be sampled within the finite intervals $[i_{k}+1,i_{k+1}]$, ensuring the well-defined nature of this process (see [Har78]).
[…]
Figure 5. The colored half-space TASEP with packed initial condition $\eta(x)=x$; the transition rate associated to edge $(-1,1)$ is $\alpha$; for $x\geq 1$, the transition rates associated to edge $(x,x+1)$ and $(-x-1,x)$ are given by $1$.
2.2. Random walk on the type BC Hecke algebra
Consider the group $B_{n}$, which is the hyperoctahedral group of signed permutations (referred to as permutations) on $n$ letters. In other words, $B_{n}$ consists of permutations $\pi$ of ${-n,\cdots,-1,1,\cdots,n}$ such that $\pi(-i)=-\pi(i)$. We treat this group as a Coxeter group, having generators $s_{0}=(-1,1)$ and $s_{k}=(k,k+1)$ for $0<k<n$. We represent elements in $B_{n}$ by the one-line notation, which means a number $i$ is placed at position $j$ if $\pi(j)=i$. The length $l(\pi)$ is defined as the minimum number of generators required to express $\pi$. We can express $l(\pi)$ as the sum $l_{0}(\pi)+l_{1}(\pi)$, where $l_{0}$ counts the occurrences of $s_{0}$ and $l_{1}$ counts the occurrences of $s_{k}$ for $k>0$ in any decomposition $\pi=s_{k_{1}}\cdots s_{k_{n}}$ with $n$ being minimal. This decomposition is well defined (see [BB05]).
We define the BC Hecke algebra to be the associative algebra $\mathcal{H}=\mathcal{H}(B_{n})$ with a linear basis $\{T_{\omega}\}_{\omega\in B_{n}}$ and multiplication which satisfy the following rules for any Coxeter generators $s_{k}$ and $\omega\in B_{n}$:
(2.1) $T_{s_{k}}T_{\omega}$ =$=\begin{cases}T_{s_{k}\omega}&\text{ if }l(s_{k}\omega)=l(\omega)+1,\\ T_{\omega}&\text{ if }l(s_{k}\omega)=l(\omega)-1.\end{cases}$
We refer to [Hum90] for the more general definition of type BC Hecke algebra, and [Buf20, HS23] for their probabilistic application in interacting particle system.
Now we consider the following random walk on the type BC Hecke algebra $\mathcal{H}(B_{n})$. We place rate 1 exponential clocks at each edge $(k,k+1)$ (associated to $s_{k}$, $k\geq 1$), and a rate $\alpha$ exponential clock at the edge $(-1,1)$ (associated to $s_{0}$). For any time $t\geq 0$, we define the stochastic process $W(t)$ which takes values in $\mathcal{H}(B_{n})$. Its initial value is $W(0)=T_{id}$ (the basis vector corresponding to the identity permutation). When one of the clock associated to $s_{k}$ rings at a certain time $\tau\in\mathbb{R}_{\geq 0}$, then we set $W(\tau)=T_{s_{k}}W(\tau^{-})$.
Let us make the link here between the random walk $W(t)$ and the colored half-space TASEP as constructed in Section 2.1. Note that the process $\{W(t),t\geq 0\}$, which takes values in the Hecke algebra $\mathcal{H}(B_{n})$, immediately induces a stochastic process $\{\omega_{t},t\geq 0\}$ on $B_{n}$. The definition of $W(t)$ and the multiplication rules 2.1 imply that the process $\{\omega_{t},t\geq 0\}$ on $B_{n}$ is exactly the colored half-space TASEP as constructed in Section 2.1: the first rule of 2.1 says that if $\omega(k)<\omega(k+1)$ and the exponential clock associated to $s_{k}$ rings, $\omega$ gets updated as $s_{k}\omega$; the second rule of 2.1 says that if $\omega(k)>\omega(k+1)$, $\omega$ remains unchanged. Note that by definition $W(0)=T_{id}$, which indicates $\omega_{0}=id$.
The reason we describe the evolution of colored half-space TASEP as the random walk on the type BC Hecke algebra is that we wish to use a key property of the Hecke algebra $\mathcal{H}(B_{n})$: the anti-involution $\iota$, which sends $\iota(T_{w})=T_{w^{-1}}$ for all $w\in B_{n}$. In more detail, for any $T_{\omega_{1}},\cdots,T_{\omega_{k}}\in\mathcal{H}(B_{n})$ we have
(2.2) $\iota(T_{\omega_{n}}T_{\omega_{n-1}}\cdots T_{\omega_{2}}T_{\omega_{1}})=\iota(T_{\omega_{1}})\iota(T_{\omega_{2}})\cdots\iota(T_{\omega_{n-1}})\iota(T_{\omega_{n}}),$
which can be straightforwardly proved by induciton in $l(\omega)$ with the use of (2.1). 2.2 was first introduced in probabilistic setting by [Buf20]. Furthermore, we use the fact for any Coxeter generator $\iota(T_{s_{k}})=T_{s_{k}}$ to induce the following lemma:
Lemma 2.1.
For any $T_{s_{k_{1}}},\cdots,T_{s_{k_{n}}}\in\mathcal{H}(B_{n})$ we have
(2.3) $\iota(T_{s_{k_{n}}}T_{s_{k_{n-1}}}\cdots T_{s_{k_{2}}}T_{s_{k_{1}}})=T_{s_{k_{1}}}T_{s_{k_{2}}}\cdots T_{s_{k_{n-1}}}T_{s_{k_{n}}}.$
Remark 2.2.
In the probabilistic setting of type A Hecke algebra, Lemma 2.1 was proved in [AHR09, Lemma 2.1], see also [BF22, Theorem 2.1]. The ASEP generalizations of [AHR09, Lemma 2.1] can be found in [AAV11, Lemma 3.1] and [BB21, Theorem 2.2].
By the time-reflection symmetry of a homogeneous Poisson process, we have the following lemma as the corollary of Lemma 2.1:
Lemma 2.3.
[…]
Remark 2.4.
See [HS23, Lemma 3.4] for the ASEP generalizations of Lemma 2.3. In the setting of type A Hecke algebra, Lemma 2.3 was also used in [AAV11, BB21, BF22, BN22].
2.3. Color-position symmetry theorem
Given a sequence of Coxeter generators denoted as $s_{k_{1}},\cdots,s_{k_{n}}$, we will explore two processes associated with this data within the framework of colored half-space TAEP.
In the first process, we commence with the configuration $T_{id}$ and sequentially apply updates $T_{s_{k_{1}}},\cdots,T_{s_{k_{n}}}$ (with $T_{s_{k_{1}}}$ being the first update). Subsequently, the continuous-time process $W(t)$ (a random walk described in section 2.2) is initiated, and the random configuration obtained at time $t$ is denoted as $\eta_{t}^{gen;1}$.
In the second process, we once again start with the initial configuration $T_{id}$ but precede it with the continuous-time process running for a duration of time $t$. Following this, we apply updates $T_{s_{k_{1}}},\cdots,T_{s_{k_{n}}}$ to the resulting random configuration in the reverse order, i.e., $T_{s_{k_{n}}},\cdots,T_{s_{k_{1}}}$. The configuration obtained is denoted as $\eta_{t}^{gen;2}$.
Additionally, for our color-position symmetry theorem, we introduce another notation $\iota(\eta^{gen;1}_{t})(x)$, representing the position of the particle with color $x$ in $\eta_{t}^{gen;1}$ (recall that $\eta_{t}^{gen;1}(x)$ denotes the color of the particle located at $x$ in the first process). Similarly, we define the notation $\iota(\eta^{gen;2}_{t})(x)$ for the second process $\eta_{t}^{gen;2}$.
Theorem 2.5.
The random configurations $\eta_{t}^{gen;1}$ and $\iota(\eta_{t}^{gen;2})$ have the same distribution, and also $\iota(\eta_{t}^{gen;1})$ have the same distribution with $\eta_{t}^{gen;2}$.
Proof.
We only prove the first part of above theorem, the proof of the second part is similar. By definition, we can write down the formula of $\eta_{t}^{gen;1}$:
$\eta_{t}^{gen;1}=W(t)T_{s_{k_{n}}}\cdots T_{s_{k_{1}}}T_{id},$
which is equal to the following formula:
$\iota(\eta_{t}^{gen;2})=\iota(T_{s_{k_{1}}}\cdots T_{s_{k_{n}}}W(t)T_{id}),$
where we use Lemma 2.1 and Lemma 2.3. ∎
Remark 2.6.
In the probabilistic setting of type A Hecke algebra, Theorem 2.5 was proved in [AAV11, Theorem 1.4], [BB21, Theorem 3.1] and [BF22, Proposition 2.3].
3. Exact distribution of the second class particles in the half-space TASEP with shocks
In this section, we use the color-position symmetry theorem introduced in Section 2 to derive the exact distribution of the second class particles in different kinds of multi-color half-line open TASEP. We focus on the one-shock case from Section 3.1 to Section 3.3, and discuss the two-shock case in Section 3.4.
3.1. Second class particles in one shock
Let $M_{1},M_{2}$ be positive integers, and consider $\pi_{M_{1},M_{2}}$ be a signed permutation of the set $\{-M_{1}-M_{2}-1,\cdots,-1,1,\cdots,M_{1}+M_{2}+1\}$ such that $\pi_{M_{1},M_{2}}(i)=M_{1}+M_{2}+2-i$ and $\pi(-i)=-\pi(i)$ for $i=1,2,\cdots,M_{1}+M_{2}+1$. We assume the length $l(\pi_{M_{1},M_{2}})=m$ and let $\pi_{M_{1},M_{2}}=s_{k_{m}}s_{k_{m-1}}\cdots s_{k_{1}}$ be a minimal length dedomposition of $\pi_{M_{1},M_{2}}$ into the Coxeter generators (there can be many such decompositions, we only choose one of them).
Consider the first process $\eta_{t}^{gen;1}$ introduced in Section 2.3 involving data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$. In this process, we initiate with the configuration $T_{id}$ and sequentially apply updates $T_{s_{k_{1}}},T_{s_{k_{2}}},\cdots,T_{s_{m}}$ to it. This can be expressed as $\eta_{0}^{gen;1}=T_{s_{k_{m}}}T_{s_{k_{m-1}}}\cdots T_{s_{k_{1}}}T_{id}$. To elaborate, within the packed initial configuration, we individually arrange integers within the ranges $[1,\cdots,M_{1}+M_{2}+1]$ and $[-M_{1}-M_{2}-1,\cdots,-1]$ in reverse order. Subsequently, we execute continuous time dynamics to attain the random configuration $\eta_{t}^{gen;1}$ up to time $t$.
In this section, we will consider two kinds of tricolor half-space TASEP which can be regarded as different projections of $\eta_{t}^{gen;1}$. Firstly, we consider a tricolor half-space TASEP denoted as $\eta_{t}$ with the initial condition
(3.1) $\eta_{0}(x)$ =$=\begin{cases}1,&\text{ if }x\leq-1\ \text{or}\ M_{1}+2\leq x\leq M_{1}+M_{2}+1,\\ 2,&\text{ if }x=M_{1}+1,\\ +\infty,&\text{otherwise}.\end{cases}$
The interpretation of $\eta_{t}$ in terms of $\eta_{t}^{gen;1}$ (with Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$) is the following. Since $\pi_{M_{1},M_{2}}(M_{1}+1)=M_{2}+1$, we identify the particle with color $M_{2}+1$ as the second class particle (denote by 2). Furthermore, particles with colors less than $M_{2}+1$ are called first class particles (denote by 1), and particles with colors greater than $M_{2}+1$ are holes (denote by +$\infty$+ ∞). Secondly, we consider a tricolor half-space TASEP denoted as $\tilde{\eta}_{t}$ with the following initial condition
(3.2) $\tilde{\eta}_{0}(x)$ =$=\begin{cases}1,&\text{ if }x\leq-1,\\ 2,&\text{ if }M_{1}+1\leq x\leq M_{1}+M_{2}+1,\\ +\infty,&\text{otherwise}.\end{cases}$
Similar to $\eta_{t}$, we can interpret $\tilde{\eta}_{t}$ in terms of $\eta_{t}^{gen;1}$ (with Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$) as following: we identify the particles with negative colors as the first class particles, the particles with colors between $[1,M_{2}+1]$ as second class particles, and the particles with colors greater than $M_{2}+1$ as holes.
Remark 3.1.
Because the dynamics on $\mathbb{N}$ and $-\mathbb{N}$ are symmetric, and the initial configurations (3.1) and (3.2) are filled with first class particles, therefore we can identity half-space TASEP $\eta_{t}$ (resp. half-space TASEP $\tilde{\eta}_{t}$) with initial condition (3.1) (resp. initial condition (3.2)) with the half-line open TASEP in Figure 1 Upper (resp. Figure 1 Bottom). That’s why we don’t change the notations. The reason is the same for the two-shock case in Section 3.4, we will not state it again later.
Both of above tricolor half-space TASEPs illustrate the second class particles in one shock. We aim to investigate the characteristics of the sole second-class particle in the half-space by examining the first process $\eta_{t}$. Simultaneously, we will explore the behavior of the multi-colored height functions in the half-space through the analysis of the second process $\tilde{\eta}_{t}$. Let $f(t)$ represent the position of the unique second class particle in the process $\eta_{t}$. Additionally, let $\mathbf{N}_{c}(x,t)$ denote the number of particles of color $c$ in the process $\tilde{\eta}_{t}$ at time $t$ that are located weakly to the right of $x\in\mathbb{R}_{\geq 0}$, where $c=1,2$, namely,
(3.3) $\mathbf{N}_{c}(x,t)=\sum_{z\geq x}\delta_{\tilde{\eta}_{t}(z),c},\quad\text{for}\ c=1,2.$
These quantities will be studied by establishing connections to a more straightforward process using the color-position symmetry.
3.2. Standard half-line open TASEP with step initial condition
In this section, we introduce a two-color half-space TASEP $\eta_{t}^{step}(x)$ starting from the step initial condition
(3.4) $\eta_{0}^{step}(x)$ =$=\begin{cases}1,&\text{ if }x\leq-1,\\ +\infty,&\text{ if }x\geq 1.\end{cases}$
We can interpret $\eta_{t}^{step}(x)$ as the projection of the colored half-space TASEP introduced in Section 2.1. In this context, if a color is $\leq-1$, it signifies the presence of a particle (denoted by 1), whereas a color $\geq 1$ indicates the existence of a hole (denoted by +$\infty$+ ∞). One can identity the two-color half-space TASEP $\eta_{t}^{step}(x)$ with the standard half-line open TASEP in Figure 6. For any $x\in\mathbb{R}_{\geq 0}$, we use $\mathcal{N}(x,t)$ to represent the number of particles weakly positioned to the right of $x$ in $\eta_{t}^{step}$, namely,
(3.5) $\mathcal{N}(x,t)=\sum_{z\geq x}\delta_{\eta_{t}^{step}(z),1}.$
[…]
Figure 6. The standard half-line open TASEP with step initial condition.
3.3. Distribution identity
In this section, we study the quantities $f(t)$, $\mathbf{N}_{1}(x,t)$ and $\mathbf{N}_{2}(x,t)$ by relating them to $\mathcal{N}(x,t)$ via the color-position symmetry theorem 2.5.
Theorem 3.2.
The probability of the unique second class particle disappearing from the half-space TASEP $\eta_{t}$ is given by:
(3.6) $\mathbb{P}(f(t)<0)=\mathbb{P}(\mathcal{N}(1,t)-\mathcal{N}(M_{1}+M_{2}+2,t)\geq M_{1}+1).$
Equivalently, the probability that the second class particle exists in the system is equal to
(3.7) $\mathbb{P}(f(t)>0)=\mathbb{P}(\mathcal{N}(1,t)-\mathcal{N}(M_{1}+M_{2}+2,t)\leq M_{1}).$
Proof.
We use Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$ introduced in Section 3.1 in the following. By the relation of $\eta_{t}$ and $\eta_{t}^{gen;1}$, we have
$\mathbb{P}(f(t)<0)=\mathbb{P}(\iota(\eta_{t}^{gen;1})(M_{2}+1)<0);$
Using the color-position symmetry theorem 2.5, we have
$\mathbb{P}((\iota\eta)_{t}^{gen;1}(M_{2}+1)<0)=\mathbb{P}(\eta_{t}^{gen;2}(M_{2}+1)<0).$
Considering the definition of $\eta_{t}^{gen;2}$ and the projection introduced in Section 3.2, the event that the color $\eta_{t}^{gen;2}(M_{2}+1)$ in the position $M_{2}+1$ of $\eta_{t}^{gen;2}$ is less than 0 is equivalent to the event in half-line TASEP (3.4) that the number of particles in the set $\{1,\cdots,M_{1}+M_{2}+1\}$ is at least $M_{1}+1$. Namely, we have the equation 3.6. ∎
Theorem 3.3.
We have the following distribution identity among $\mathbf{N}_{1}(1,t)$, $\mathbf{N}_{2}(1,t)$ and $\mathcal{N}(x,t)$:
(3.8) $\left(\mathbf{N}_{1}(1,t),\mathbf{N}_{2}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\mathcal{N}(1,t),\min\left\{M_{1}+M_{2}+1-\mathcal{N}(1,t)+\mathcal{N}(M_{1}+M_{2}+2,t),M_{2}+1\right\}\right),$
where by =$\stackrel{{\scriptstyle d}}{{=}}$ we denote the equality in distribution.
Proof.
We use Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$ introduced in Section 3.1 in the folloing. By the relation of $\tilde{\eta}_{t}$ and $\eta_{t}^{gen;1}$, we have
$\left(\mathbf{N}_{1}(1,t),\mathbf{N}_{2}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\#\{i:i<0,\iota(\eta_{t}^{gen;1})(i)\geq 1\},\#\{i:1\leq i\leq M_{2}+1,\iota(\eta_{t}^{gen;1})(i)\geq 1\}\right),$
Using the color-position symmetry theorem 2.5, we have
(#$\left(\#\{i:i<0,\iota(\eta_{t}^{gen;1})(i)\geq 1\},\#\{i:1\leq i\leq M_{2}+1,\iota(\eta_{t}^{gen;1})(i)\geq 1\}\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\},\#\{i:1\leq i\leq M_{2}+1,\eta_{t}^{gen;2}(i)\geq 1\}\right),$
Similar with the proof of Theorem 3.2, we can get 3.11. ∎
Remark 3.4.
Theorem 3.2 and Theorem 3.3 are similar to the results in [BB21, Proposition 6.1] and [BF22, Proposition 2.5], but in the half-space geometry. Different from full-space geometry, we lose the property of translation-invariance. That is why [BF22, Proposition 2.5] can give an accurate description of the distribution function, while we only get the probability distribution around the origin.
Remark 3.5.
We can generalize the results in Theorem 3.2 and Theorem 3.3 to the distribution identities of distribution functions by introducing the half-line open TASEP with the shifted step initial condition(denoted by $\eta_{t}^{step(y)}$, $y\in\mathbb{N}\cup-\mathbb{N}$) :
(3.9) $\eta_{0}^{step(y)}(x)$ =$=\begin{cases}1,&\text{ if }x\leq y,\\ +\infty,&\text{ if }x>y.\end{cases}$
where $x\in\mathbb{N}\cup-\mathbb{N}$. We can interpret $\eta_{t}^{step(y)}$ as the projection of the colored half-space TASEP introduced in Section 2.1: if a color is $\leq y$, we identify it as a particle (denoted by 1), whereas a color $>y$ is regarded as a hole (denoted by +$\infty$+ ∞). For any $x\in\mathbb{R}$, we use $\mathcal{N}^{y}(x,t)$ to represent the number of particles which are weakly to the right of $x$ in $\eta_{t}^{step(y)}$, namely,
(3.10) $\mathcal{N}^{y}(x,t)=\sum_{z\geq x}\delta_{\eta_{t}^{step(y)}(z),1}.$
Then the equation (3.6) can be generalized to the following form:
$\mathbb{P}(f(t)\leq y)=\mathbb{P}(\mathcal{N}^{y}(1,t)-\mathcal{N}^{y}(M_{1}+M_{2}+2,t)\geq M_{1}+1).$
And the equation (3.11) can be generalized to the following form:
(3.11) $\left(\mathbf{N}_{1}(y,t),\mathbf{N}_{2}(y,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\mathcal{N}^{y}(1,t)-y,\min\left\{M_{1}+M_{2}+1-\mathcal{N}^{y}(1,t)+\mathcal{N}^{y}(M_{1}+M_{2}+2,t),M_{2}+1\right\}\right).$
Since we know very little about the asymptotic properties of $\mathcal{N}^{y}(x,t)$, this method is far to reach.
Remark 3.6.
One can also consider generalizing the results in Theorem 3.2 and Theorem 3.3 to the distribution identities of distribution functions by introducing the shift-invariance of colored half-space TASEP. Let $h^{\leq i}(x,t)$ denote the number of particles of color $\leq i$ which are weakly to the right of $x$. Then we have
$\mathbb{P}(f(t)\leq y)=\mathbb{P}(h^{\leq y}(1,t)-h^{\leq y}(M_{1}+M_{2}+1,t)\geq M_{1}+1).$
We need to derive the shift-invariance theorem to set up the connection between $h^{\leq y}(x,t)$ and $h^{\leq-1}(1,t)$, for any $x,y\in\mathbb{N}\cup-\mathbb{N}$. We have noticed that there exists specific shift-invariance theorem in [He22], but it doesn’t fit our situation.
3.4. Second class particles in two shocks
In this section, we consider two more multi-color half-space TASEPs which have two-shock initial conditions. Let $N,M$ be positive integers. The first multi-color half-space TASEP has three colors, we denote it as $\xi_{t}$ and set the following initial condition:
(3.12) $\xi_{0}(x)$ =$=\begin{cases}1,&\text{ if }x\leq-1\ \text{or}\ \ N+1\leq x\leq N+M,\\ 2,&\text{ if }x=N+2M+1,\\ 1,&\text{ if }N+2M+2\leq x\leq 2N+2M+1,\\ +\infty,&\text{otherwise}.\end{cases}$
The second multi-color half-space TASEP has four colors, we denote it as $\tilde{\xi}_{t}$ and assign the following initial condition:
(3.13) $\tilde{\xi}_{0}(x)$ =$=\begin{cases}1,&\text{ if }x\leq-1,\\ 2,&\text{ if }N+1\leq x\leq N+M,\\ 3,&\text{ if }2M+N+1\leq x\leq 2M+2N+1,\\ +\infty,&\text{otherwise}.\end{cases}$
We denote the position of the unique second class particle in the process $\xi_{t}$ by $g(t)$. Further, let $\mathsf{N}_{i}(x,t)$ be the number of particles of color $i$ in $\tilde{\xi}_{t}$ which are weakly to the right of $x\in\mathbb{R}$, where $i\in\{1,2,3\}$.
Theorem 3.7.
The probability of the unique second class particle disappearing from the half-space TASEP $\xi_{t}$ is given by:
(3.14) $\mathbb{P}\left(g(t)<0\right)=\mathbb{P}\left(\mathcal{N}(N+M+1,t)-\mathcal{N}(2N+2M+2,t)\right.\\ \left.+\min\{N,\mathcal{N}(1,t)-\mathcal{N}(N+M+1,t)\}\geq N+M+1\right);$
And we also have the following distribution identity among $\mathsf{N}_{1}(1,t)$, $\mathsf{N}_{2}(1,t)$, $\mathsf{N}_{3}(1,t)$ and $\mathcal{N}(x,t)$:
(3.15) $\left(\mathsf{N}_{1}(1,t),\mathsf{N}_{2}(1,t),\mathsf{N}_{3}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\bigl{(}\mathcal{N}(1,t),\min\{M,M+N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\},\bigr{.}\\ \bigl{.}\min\bigl{\{}N+1,M+N+1-\mathcal{N}(M+N+1,t)+\mathcal{N}(2M+2N+2,t)\bigr{.}\bigr{.}\\ \bigl{.}\bigl{.}+\max\{0,N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\bigr{\}}\bigr{)}.$
Proof.
Let $\pi_{1}$ be a signed permutation of the set $\{-2M-2N-1,\cdots,-M-1,M+1,\cdots,2M+2N+1\}$ such that $\pi_{1}(M+1+i)=\pi_{1}(2M+2N+1-i)$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+2N$. And let $\pi_{2}$ be a signed permutation of the set $\{-M-N,\cdots,-1,1,\cdots,M+N\}$ such that $\pi_{2}(1+i)=\pi_{1}(M+N-i)$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+N-1$. Consider a minimal length decomposition of permutation $\pi_{2}\cdot\pi_{1}$ into Coxeter generators: $\pi_{2}\cdot\pi_{1}=s_{\ell_{p}}s_{\ell_{p-1}}\cdots s_{\ell_{1}}$, where $l(\pi_{2}\cdot\pi_{1})=p$.
We use the data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct the first process $\eta_{t}^{gen;1}$ introduced in Section 2.3. We can get $\xi_{t}$ from $\eta_{t}^{gen;1}$(with Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$) by identifying the particle with color $M+N+1$ as the second class particle (denote by 2), particles with colors less than $M+N+1$ as first class particles (denote by 1), and particles with colors greater than $M+N+1$ are holes (denote by +$\infty$+ ∞). That means:
$\mathbb{P}(g(t)<0)=\mathbb{P}(\iota(\eta_{t}^{gen;1})(M+N+1)<0).$
Additionally, $\tilde{\xi}_{t}$ can be interpreted as the projection of $\eta_{t}^{gen;1}$(with Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$) by the following way: we identify the particles with negative colors as first class particles, colors between $[1,M]$ as second class particles, colors between $[M+1,M+N+1]$ as third class particles (denote by 3), and colors greater than $M+N+1$ as holes. This gives:
$\left(\mathsf{N}_{1}(1,t),\mathsf{N}_{2}(1,t),\mathsf{N}_{3}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\left(\#\{i:i<0,\iota(\eta_{t}^{gen;1})(i)\geq 1\},\right.\\ \left.\#\{i:1\leq i\leq M,\iota(\eta_{t}^{gen;1})(i)\geq 1\},\#\{i:M+1\leq i\leq M+N+1,\iota(\eta_{t}^{gen;1})(i)\geq 1\}\right).$
The color-position symmetry theorem 2.5 implies:
$\mathbb{P}(g(t)<0)=\mathbb{P}(\eta_{t}^{gen;2}(M+N+1)<0),$
and the following relation:
$\left(\mathsf{N}_{1}(1,t),\mathsf{N}_{2}(1,t),\mathsf{N}_{3}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\left(\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\},\right.\\ \left.\#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\},\#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\right),$
where we use Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct $\eta_{t}^{gen;2}$(recall the definition introduced in Section 2.3). Considering the definition of $\eta_{t}^{gen;2}$ and the projection introduced in Section 3.2, the color of the particle in position $M+N+1$ is less than 0 in $\eta_{t}^{gen;2}$ is equivalent to the event in half-line TASEP 3.4 that the number of particles in the set $\{M+1,\cdots,2M+2N+1\}$ is at least $M+N+1$. Namely, we have the equation 3.14. Similarly, we have the following distribution identity:
#$\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \#\{i:i>0,\eta_{t}^{gen;2}(i)\leq-1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \mathcal{N}(1,t),\\ \#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \min\{M,M+N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\\ \#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\stackrel{{\scriptstyle d}}{{=}}\min\bigl{\{}N+1,M+N+1-\mathcal{N}(M+N+1,t)\bigr{.}\\ \bigl{.}+\mathcal{N}(2M+2N+2,t)+\max\{0,N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\bigr{\}},$
which lead to the proof of 3.15. ∎
4. Asymptotic distribution of second class particles under constant scaling
In this section, we fix the size of the shocks as constants, and consider the asymptotic distribution of the second class particles introduced in Section 3. In Section 4.1, we review the stationary measure of the standard half-line open TASEP with the step initial condition, and derive a combinatorial formula for the stationary measure at $\alpha>\frac{1}{2}$. In Section 4.2 we give the exact formulas of the asymptotic behavior of the second class particles.
4.1. Stationary measure of half-line open TASEP
According to Liggett’s results on half-line open ASEP [Ligg75], we can characterize the stationary measure of half-line TASEP with step initial condition (denoted by $\eta_{t}^{step}$) as follows:
Theorem 4.1.
Let parameter $\alpha\in(0,1)$ corresponds to the density of particles that the reservoir imposes at site 1. When $\alpha\leq\frac{1}{2}$, the process $\eta_{t}^{step}$ admits stationary measures that are product i.i.d. Bernoulli with parameter $\alpha$. When $\alpha>\frac{1}{2}$, stationary measures (denoted by $\mu_{1/2}^{\alpha}$) of process $\eta_{t}^{step}$ are spatially correlated and more complicated: there is a rarefaction fan with a density of particles $\alpha$ near the origin and density $\frac{1}{2}$ at +$\infty$+ ∞.
Following the seminal work of [DEHP93], Grosskinsky [Gro04] has given a explicit characterisation of the stationary measure $\mu_{1/2}^{\alpha}$ of half-line open TASEP with a matrix product ansatz:
Theorem 4.2.
The $L$ dimensional distributions of stationary measures $\mu_{1/2}^{\alpha}$ for the process $\eta_{t}^{step}$ are given by the following matrix product ansatz:
(4.1) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{L}=\eta_{L}\}=\left\langle w\right|\prod_{x=1}^{L}\left(\eta_{x}D+(1-\eta_{x})E\right)\left|v\right\rangle,$
where $L\in\mathbb{N}$, $(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, the nonnegative matrices $D$, $E$ and vectors $\left\langle w\right|$ and $\left|v\right\rangle$ satisfy the following algebraic relations with $c=\frac{1}{4}$:
(4.2) $DE=c(D+E),$
(4.3) $\alpha\left\langle w\right|E=c\left\langle w\right|,$
(4.4) $(D+E)\left|v\right\rangle=\left|v\right\rangle,$
(4.5) $\left\langle w\right|\left|v\right\rangle=1.$
With rules (4.2), (4.3) and (4.5), the probability of any cylinder set $(\eta_{1},\cdots,\eta_{L})$ can be directly written as a linear combination of $\left\langle w\right|D^{k}\left|v\right\rangle$:
$\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{L}=\eta_{L}\}=\sum_{k=0}^{\ell(\eta)}d_{k}(\eta)\left\langle w\right|D^{k}\left|v\right\rangle,$
where $\ell(\eta)=\#\{i:\eta_{i}=1\}$. Using (4.4), Grosskinsky [Gro04] has noticed that $\left\langle w\right|D^{k}\left|v\right\rangle$ satisfies the following recursion:
$\left\langle w\right|D^{k+2}\left|v\right\rangle=\left\langle w\right|D^{k+1}(D+E)\left|v\right\rangle-\left\langle w\right|D^{k}DE\left|v\right\rangle=\left\langle w\right|D^{k+1}\left|v\right\rangle-c\left\langle w\right|D^{k}\left|v\right\rangle,$
with initial values $\left\langle w\right|D^{0}\left|v\right\rangle=1$ and $\left\langle w\right|D\left|v\right\rangle=1-c/\alpha$, and the solutions of this linear recursion are given by the following formulas:
(4.6) $\left\langle w\right|D^{k}\left|v\right\rangle=\frac{1}{2^{k}}\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),\quad 0\leq\ell(\eta)\leq k.$
In this section, we are focused on the combinatorial expression of the coefficients $d_{k}(\eta)$ for any $\eta$ and $0\leq k\leq\ell(\eta)$.
We suppose that $\eta=(\eta_{1},\cdots,\eta_{L})$ consists of $n$ clusters of $(1,\cdots,1)$ with length $\sigma_{i}$ and $n+1$ clusters of $(0,\cdots,0)$ with length $\tau_{j}$, which has the following form:
(4.7) $\eta=(\underbrace{0,\cdots,0}_{\tau_{0}},\underbrace{1,\cdots,1}_{\sigma_{1}},\underbrace{0,\cdots,0}_{\tau_{1}},\underbrace{1,\cdots,1}_{\sigma_{2}},\cdots,\underbrace{0,\cdots,0}_{\tau_{n-1}},\underbrace{1,\cdots,1}_{\sigma_{n}},\underbrace{0,\cdots,0}_{\tau_{n}})$
where $0\leq n\leq\lceil L/2\rceil$, $1\leq\sigma_{i}\leq L$ for $i\in\{1,2,\cdots,n\}$, $1\leq\tau_{j}\leq L$ for $j\in\{1,\cdots,n-1\}$ and $0\leq\tau_{0},\tau_{n}\leq L$. Then the expression of (4.1) is the following form:
$\mu_{\alpha}^{1/2}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{L}=\eta_{L}\}$
$=$ $\left\langle w\right|\underbrace{E\ldots E}_{\tau_{0}}\underbrace{D\ldots D}_{\sigma_{1}}\underbrace{E\ldots E}_{\tau_{1}}\underbrace{D\ldots D}_{\sigma_{2}}\cdots\underbrace{E\ldots E}_{\tau_{n-1}}\underbrace{D\ldots D}_{\sigma_{n}}\underbrace{E\ldots E}_{\tau_{n}}\left|v\right\rangle$
$=$ $(c/\alpha)^{\tau_{0}}\left\langle w\right|\underbrace{D\ldots D}_{\sigma_{1}}\underbrace{E\ldots E}_{\tau_{1}}\underbrace{D\ldots D}_{\sigma_{2}}\cdots\underbrace{E\ldots E}_{\tau_{n-1}}\underbrace{D\ldots D}_{\sigma_{n}}\underbrace{E\ldots E}_{\tau_{n}}\left|v\right\rangle$
::=:= $(c/\alpha)^{\tau_{0}}F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n}),$
where we use the algebraic relation (4.3) and the expression (4.7) of $\eta$. Our main task is to express $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ as a linear combination of $\left\langle w\right|D^{k}\left|v\right\rangle$.
As an appetizer, we first give the explicit combinatorial formula of $F_{1}(\sigma_{1},\tau_{1})$ with only one pair of blocks $\underbrace{D\ldots D}_{\sigma_{1}}\underbrace{E\ldots E}_{\tau_{1}}$. We achieve this goal by introducing the one-layer DEHP-tree.
Definition 4.3.
The one-layer DEHP-tree of $(\sigma_{1},\tau_{1})$ is a rooted tree with DEHP weight, which is defined by the following way:
  • start from a node represented by a tuple $(\sigma_{1},\tau_{1})$, which we call the root;
[…]
  • any node $(0,y)$ with $y\geq 1$ has only one child: the left child $(0,y-1)$, the parent $(0,y)$ is connected with the left child $(0,y-1)$ by a path weighted by $c/\alpha$;
  • end with the nodes $(x,0)$ where $x=0,1,\cdots,\sigma_{1}$, which we call the endpoints.
Note that there exists at least one directed path from the root $(\sigma_{1},\tau_{1})$ to any child $(x,y)$, we define its weight as the product of weights on the path. Define the partition function of a node $(x,y)$ as the sum of weights of all possible paths from the root $(\sigma_{1},\tau_{1})$ to that node, we denote it by $Z_{x,y}(\sigma_{1},\tau_{1})$. In particular, we denote the partition function of endpoints $(k,0)$ as $Z_{k}(\sigma_{1},\tau_{1})$. (see Figure 7 for example).
$(1,2)$$c$$(1,1)$$c$$(0,2)$$c$$(1,0)$$c$$(0,1)$$c/\alpha$$(0,0)$$c/\alpha$
Figure 7. The one-layer DEHP-tree of $DEE$.
Proposition 4.4.
We can give the explicit combinatorial formula of $F(\sigma_{1},\tau_{1})$ by the above one-layer DEHP-tree:
(4.8) $F(\sigma_{1},\tau_{1})=\sum_{k=0}^{\sigma_{1}}Z_{k}(\sigma_{1},\tau_{1})\left\langle w\right|D^{k}\left|v\right\rangle.$
Proof.
By the result (4.6), we conclude the proof by identifing the following relation:
$(x,y)\iff\underbrace{D\ldots D}_{x}\underbrace{E\ldots E}_{y}$
Now we are in a position to extend the definition of one-layer DEHP-tree to $n$-layer DEHP-tree, and then give the explicit combinatorial formula of $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ with arbitrary $n$ pairs of blocks $\prod_{i=1}^{n}\underbrace{D\ldots D}_{\sigma_{i}}\underbrace{E\ldots E}_{\tau_{i}}$. We introduce the notation $\tilde{\Psi}_{i}=\sum_{j=1}^{i}\sigma_{j}$ and $\tilde{\Phi}_{i}=\sum_{j=1}^{i}\tau_{j}$.
Definition 4.5.
The $n$-layer DEHP-tree of $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$ is a rooted tree with DEHP weight, which is defined by the following recursive way:
  • start from an one-layer DEHP-tree of $(\tilde{\Psi}_{1},\tilde{\Phi}_{1})$;
  • construct the $i$-layer DEHP-tree of $(\tilde{\Psi}_{i},\tilde{\Phi}_{i})$ from the $(i-1)$-layer DEHP-tree of $(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})$: we add $(\sigma_{i},\tau_{i})$ to every node of the $(i-1)$-layer DEHP-tree of $(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})$ and then continue the tree construction by performing the generation of one-layer DEHP-tree of $(\tilde{\Psi}_{i},\tau_{i})$.
  • end with the endpoints $(x,0)$ where $x=0,1,\cdots,\tilde{\Psi}_{n}$.
Define the partition function of a node $(x,y)$ as the sum of weights of all possible paths from the root $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$ to that node, we denote it by $Z_{x,y}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$. In particular, we denote the partition function of endpoints $(k,0)$ as $Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$. (see Figure 8 for examples).
Proposition 4.6.
We can give the explicit formula of $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ by the above $n$-layer DEHP-tree:
(4.9) $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})=\sum_{k=0}^{\tilde{\Psi}_{n}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})\left\langle w\right|D^{k}\left|v\right\rangle.$
Proof.
We can conclude the proof from the definition 4.5. ∎
$(3,3)$$c$$(3,2)$$c$$(2,3)$$c$$(3,1)$$c$$(2,2)$$c$$(3,0)$$c$$(2,1)$$(2,0)$$c$$c$$(1,1)$$c$$c$$(1,0)$$(0,1)$$(0,0)$$c/\alpha$$c/\alpha$$c/\alpha$$(5,5)$$c$$(5,4)$$c$$(4,5)$$c$$(5,3)$$c$$(4,4)$$c$$(5,2)$$c$$(4,3)$$c$$(5,1)$$c$$(4,2)$$c$$c$$(3,3)$$c$$c$$c$$c$$c$$c$$(5,0)$$(4,1)$$(3,2)$$(2,3)$$c$$c$$c$$c$$(4,0)$$(3,1)$$(2,2)$$c$$c$$c$$c$$(3,0)$$(2,1)$$(1,2)$$c$$c$$c$$c$$(2,0)$$(1,1)$$(0,2)$$c$$c$$(1,0)$$(0,1)$$c/\alpha$$c/\alpha$$c/\alpha$$c/\alpha$$c/\alpha$$(0,0)$
Figure 8. Left panel: The two-layer DEHP-tree of $DEEDDE$; Right panel: The three-layer DEHP-tree of $DEEDDEDDEE$.
Corollary 4.7.
For any $\eta$ with the form 4.7, the stationary measure 4.1 has the explicit formula:
(4.10) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\tilde{\Psi}_{n}}\frac{c^{\tau_{0}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})}{2^{k}\alpha^{\tau_{0}}}\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),$
where $c=\frac{1}{4}$ and $Z_{k}(\Psi_{n},\Phi_{n})$ is given by the partition function of $(k,0)$ in the $n$-layer DEHP-tree.
Remark 4.8.
One can check the construction of DEHP-tree in Definition 1.1 is equivalent to the above construction. The difference is that we incorporate the term $\left\langle w\right|\underbrace{E\ldots E}_{\tau_{0}}$ into the Definition 1.1. The reason we do this is that the expression of formula is more neat:
(4.11) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
where $\Psi_{1}(\eta)$ and $Z_{k}(\eta)$ are defined in Section 1.3.
4.2. Explicit formulas for asymptotic distribution of the second class particles
Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
Theorem 4.9.
The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
(4.12) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.13) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
Let $0\leq m\leq M_{2}$, the asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
(4.14) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.15) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ have the following relationships:
(4.16) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=M_{2}+1).$
The asymptotic distribution of the unique second class particle disappearing from $\eta_{t}$ is given by the following formula:
(4.17) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\sum_{m=0}^{M_{2}}\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m).$
Proof.
The proof directly follows from Theorem 3.2, Theorem 3.3, Theorem 4.2 and Remark 4.8. ∎
Our second main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$:
Theorem 4.10.
The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
(4.18) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.19) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
Let $0\leq s\leq N$, the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
(4.20) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(4.21) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ have the following relationships:
(4.22) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=N+1).$
The asymptotic distribution of the unique second class particle disappearing from $\xi_{t}$ is given by the following formula:
(4.23) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\sum_{s=0}^{N}\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s).$
Proof.
The proof directly follows from Theorem 3.7, Theorem 4.2 and Remark 4.8. ∎
5. Asymptotic distribution of second class particles under KPZ-type scaling
In this Section, we study the distribution of the second class particles in the shocks on the KPZ-type scaling. First, we need the known asymptotics of the standard half-line TASEP from the step initial condition.
Definition 5.1 ([Rai00] Section 8).
For a $2\times 2$-matrix valued skew-symmetric kernel,
$K(x,y)=\begin{pmatrix}\mathsf{K}_{11}(x,y)&\mathsf{K}_{12}(x,y)\\ \mathsf{K}_{21}(x,y)&\mathsf{K}_{22}(x,y)\end{pmatrix},\ \ x,y\in\mathbb{X},$
we define its Fredholm Pfaffian by the series expansion
(5.1) $\mathrm{Pf}\big{(}J+K\big{)}_{\mathbb{L}^{2}(\mathbb{X},\mu)}=1+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{\mathbb{X}}\dots\int_{\mathbb{X}}\mathrm{Pf}\Big{(}K(x_{i},x_{j})\Big{)}_{i,j=1}^{k}\mathrm{d}\mu^{\otimes k}(x_{1}\dots x_{k}),$
provided the series converges, and we recall that for an skew-symmetric $2k\times 2k$ matrix $A$, its Pfaffian is defined by
(5.2) $\mathrm{Pf}(A)=\frac{1}{2^{k}k!}\sum_{\sigma\in\mathcal{S}_{2k}}\mathrm{sign}(\sigma)a_{\sigma(1)\sigma(2)}a_{\sigma(3)\sigma(4)}\dots a_{\sigma(2k-1)\sigma(2k)}.$
The kernel $J$ is defined by
$J(x,y)=\begin{cases}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}&\text{ if }x=y,\\ 0&\text{ if }x\neq y.\end{cases}$
Definition 5.2.
We define the half-space ${Airy}_{2}$ process $\mathcal{A}_{\varpi}^{HS}$ via its finite dimensional distributions:
(5.3) $\mathbb{P}(\mathcal{A}_{\varpi}^{HS}(x_{1})\leq\eta_{1},\cdots,\mathcal{A}_{\varpi}^{HS}(x_{n})\leq\eta_{n})=\mathrm{Pf}(J-\mathsf{K}^{\rm cross})_{\mathbb{L}^{2}(\mathbb{D}_{k}(x_{1},\cdots,x_{k}))},$
where the right hand side is the Fredholm Pfaffian (see Definition 5.1) of some kernel $\mathsf{K}^{\rm cross}$ (depending on $\varpi$ and the $\eta_{i}$). The kernel $\mathsf{K}^{\rm cross}$ was introduced in [BBCS18a, Section 2.5] which can be written as
$\mathsf{K}^{\rm cross}(i,x;j,y)=\mathsf{I}^{\rm cross}(i,x;j,y)+\mathsf{R}^{\rm cross}(i,x;j,y),$
where we have
$\mathsf{I}_{11}^{\rm cross}(i,x;j,y)=$ $\frac{1}{(2\mathbf{i}\pi)^{2}}\int_{\mathcal{C}_{1}^{\pi/3}}\mathrm{d}z\int_{\mathcal{C}_{1}^{\pi/3}}\mathrm{d}w\frac{z+\eta_{i}-w-\eta_{j}}{z+w+\eta_{i}+\eta_{j}}\frac{z+\varpi+\eta_{i}}{z+\eta_{i}}\frac{w+\varpi+\eta_{j}}{w+\eta_{j}}e^{z^{3}/3+w^{3}/3-xz-yw},$
$\mathsf{I}_{12}^{\rm cross}(i,x;j,y)=$ $\frac{1}{(2\mathbf{i}\pi)^{2}}\int_{\mathcal{C}_{a_{z}}^{\pi/3}}\mathrm{d}z\int_{\mathcal{C}_{a_{w}}^{\pi/3}}\mathrm{d}w\frac{z+\eta_{i}-w+\eta_{j}}{2(z+\eta_{i})(z+\eta_{i}+w-\eta_{j})}\frac{z+\varpi+\eta_{i}}{-w+\varpi+\eta_{j}}e^{z^{3}/3+w^{3}/3-xz-yw},$
$\mathsf{I}_{21}^{\rm cross}(i,x;j,y)=$ $-\mathsf{I}_{12}^{\rm cross}(y,x),$
$\mathsf{I}_{22}^{\rm cross}(i,x;j,y)=$ $\frac{1}{(2\mathbf{i}\pi)^{2}}\int_{\mathcal{C}_{b_{z}}^{\pi/3}}\mathrm{d}z\int_{\mathcal{C}_{b_{w}}^{\pi/3}}\mathrm{d}w\frac{z-\eta_{i}-w+\eta_{j}}{4(z-\eta_{i}+w-\eta_{j})}\frac{e^{z^{3}/3+w^{3}/3-xz-yw}}{(z-\varpi-\eta_{i})(w-\varpi-\eta_{j})}.$
The contours in $\mathsf{I}_{12}^{\rm cross}$ are chosen so that $a_{z}>-\eta_{i}$, $a_{z}+a_{w}>\eta_{j}-\eta_{i}$ and $a_{w}<\varpi+\eta_{j}$. The contours in $\mathsf{I}_{22}^{\rm cross}$ are chosen so that $b_{z}>\eta_{i}$, $b_{z}>\eta_{i}+\varpi$ and $b_{w}>\eta_{j}$, $b_{w}>\eta_{j}+\varpi$. We have $\mathsf{R}_{11}^{\rm cross}(i,x;j,y)=0$, and $\mathsf{R}_{12}^{\rm cross}(i,x;j,y)=0$ when $i\geqslant j$. When $i<j$,
$\mathsf{R}_{12}^{\rm cross}(i,x;j,y)=\frac{-\exp\left(\frac{-(\eta_{i}-\eta_{j})^{4}+6(x+y)(\eta_{i}-\eta_{j})^{2}+3(x-y)^{2}}{12(\eta_{i}-\eta_{j})}\right)}{\sqrt{4\pi(\eta_{j}-\eta_{i})}},$
which may also be written as
$\mathsf{R}_{12}^{\rm cross}(i,x;j,y)=-\int_{-\infty}^{+\infty}\mathrm{d}\lambda e^{-\lambda(\eta_{i}-\eta_{j})}{\rm Ai}(x_{i}+\lambda){\rm Ai}(x_{j}+\lambda).$
The kernel $\mathsf{R}_{22}^{\rm cross}$ is antisymmetric, and when $x_{i}-\eta_{i}>y_{j}-\eta_{j}$ we have
$\mathsf{R}_{22}^{\rm cross}(i,x;j,y)=\frac{-1}{4}\frac{1}{2\mathbf{i}\pi}\int_{\mathcal{C}_{c_{z}}^{\pi/3}}\mathrm{d}z\frac{\exp\big{(}(z+\eta_{i})^{3}/3+(\varpi+\eta_{j})^{3}/3-x(z+\eta_{i})-y(\varpi+\eta_{j})\big{)}}{\varpi+z}\\ +\frac{1}{4}\frac{1}{2\mathbf{i}\pi}\int_{\mathcal{C}_{c_{z}}^{\pi/3}}\mathrm{d}z\frac{\exp\big{(}(z+\eta_{j})^{3}/3+(\varpi+\eta_{i})^{3}/3-y(z+\eta_{j})-x(\varpi+\eta_{i})\big{)}}{\varpi+z}\\ -\frac{1}{2}\frac{1}{2\mathbf{i}\pi}\int_{\mathcal{C}_{d_{z}}^{\pi/3}}\mathrm{d}z\frac{z\exp\big{(}(z+\eta_{i})^{3}/3+(-z+\eta_{j})^{3}/3-x(z+\eta_{i})-y(-z+\eta_{j})\big{)}}{(\varpi+z)(\varpi-z)},\$
where the contours are chosen so that $c_{z}<-\varpi$ and $d_{z}$ is between $-\varpi$ and $\varpi$. In the above, we use the notation $\mathcal{C}_{a}^{\varphi}$: the union of two semi-infinite rays departing $a\in\mathbb{C}$ with angles $\varphi$ and $-\varphi$, we assume that the contour is oriented from $a+\infty e^{-i\varphi}$ to $a+\infty e^{+i\varphi}$.
Theorem 5.3 ([BBCS18a, BBCS18b]).
Let $\{\mathcal{N}(x,t)\}_{x\in\mathbb{Z}_{>0}}$ be the currents of the half-space TASEP started from empty initial configuration. Assume that $u\in\mathbb{R}$. Setting $\alpha=\frac{1+2^{4/3}\varpi t^{-1/3}}{2}$, where $\varpi\in\mathbb{R}$. We have
(5.4) $\lim_{t\to\infty}\frac{\mathcal{N}(2^{\frac{1}{3}}t^{\frac{2}{3}}u,t)-\frac{t}{4}+2^{-\frac{2}{3}}t^{\frac{2}{3}}u-2^{-\frac{4}{3}}t^{\frac{1}{3}}u^{2}}{-2^{-\frac{4}{3}}t^{\frac{1}{3}}}=\mathcal{A}_{\varpi}^{HS}(u),$
where on the right hand side $\mathcal{A}_{\varpi}^{HS}(x)$ stands for the half-space ${Airy}_{2}$ process, and the convergence is in the sense of finite-dimensional distributions.
Theorem 5.4.
Assume that $M_{1}=M_{2}=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$. Setting $\alpha=\frac{1+2^{4/3}\varpi t^{-1/3}}{2}$, where $\varpi\in\mathbb{R}$. The asymptotic distribution of the unique second class particle disappearing from the half-space TASEP $\eta_{t}$ is given by:
(5.5) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\mathbb{P}\left(\mathcal{A}_{\varpi}^{HS}(0)+2^{4/3}a^{2}\leq\mathcal{A}_{\varpi}^{HS}(2^{2/3}a)\right)$
The asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
(5.6) $\lim_{t\to\infty}\frac{\mathbf{N}_{2}(1,t)-at^{2/3}}{t^{1/3}}\stackrel{{\scriptstyle d}}{{=}}\min\left\{2^{-4/3}\mathcal{A}_{\varpi}^{HS}(0)+a^{2}-2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{2/3}a),0\right\}$
Proof.
By Theorem 3.2, the asymptotic distribution of the unique second class particle disappearing in the half-space TASEP $\eta_{t}$ is given by following formula:
(5.7) $\lim_{t\to\infty}\mathbb{P}(\mathcal{N}(1,t)-\mathcal{N}(2at^{\frac{2}{3}},t)\geq at^{\frac{2}{3}}).$
By Theorem 5.3, we have
$\mathcal{N}(1,t)=\frac{t}{4}-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(0)+o(t^{\frac{1}{3}}),$
$\mathcal{N}(2at^{\frac{2}{3}},t)=\frac{t}{4}-at^{\frac{2}{3}}+a^{2}t^{\frac{1}{3}}-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)+o(t^{\frac{1}{3}}),$
where $o(t^{\frac{1}{3}})$ means that after dividing by $t^{\frac{1}{3}}$ this term will converge to 0 in probability. Therefore, in the $t\to\infty$ limit 5.7 converges to
$\mathbb{P}\left(2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2}\geq 0\right)$
then we have (5.5). Similarly, one can get (5.6) by using Theorem 3.3:
$\lim_{t\to\infty}\mathbf{N}_{2}(1,t)\stackrel{{\scriptstyle d}}{{=}}\lim_{t\to\infty}\min\left\{2at^{2/3}-\mathcal{N}(1,t)+\mathcal{N}(2at^{\frac{2}{3}},t),at^{2/3}\right\}.$
Theorem 5.5.
Assume that $M=N=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$. Setting $\alpha=\frac{1+2^{4/3}\varpi t^{-1/3}}{2}$, where $\varpi\in\mathbb{R}$. The asymptotic distribution of the unique second class particle disappearing from the half-space TASEP $\xi_{t}$ is given by:
(5.8) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\mathbb{P}\left(2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{5}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-3a^{2}\right.\\ \left.+\min\left\{2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2},0\right\}\geq 0\right).$
The asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
(5.9) $\lim_{t\to\infty}\frac{\mathsf{N}_{3}(1,t)-at^{2/3}}{t^{1/3}}\stackrel{{\scriptstyle d}}{{=}}\min\left\{0,2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{2/3}a)+3a^{2}-2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{5/3}a)\right.\\ \left.+\max\{0,2^{-4/3}\mathcal{A}_{\varpi}^{HS}(0)+a^{2}-2^{-4/3}\mathcal{A}_{\varpi}^{HS}(2^{2/3}a)\}\right\}.$
Proof.
By using (3.14), the asymptotic distribution of the unique second class particle disappearing from the half-space TASEP $\xi_{t}$ is given by:
(5.10) $\lim_{t\to\infty}\mathbb{P}(\mathcal{N}(2at^{\frac{2}{3}},t)-\mathcal{N}(4at^{\frac{2}{3}},t)+\min\{\mathcal{N}(1,t)-\mathcal{N}(2at^{\frac{2}{3}},t),at^{\frac{2}{3}}\}\geq 2at^{\frac{2}{3}})$
By Theorem 5.3, we have
$\mathcal{N}(1,t)=\frac{t}{4}-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(0)+o(t^{\frac{1}{3}}),$
$\mathcal{N}(2at^{\frac{2}{3}},t)=\frac{t}{4}-at^{\frac{2}{3}}+a^{2}t^{\frac{1}{3}}-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)+o(t^{\frac{1}{3}}),$
$\mathcal{N}(4at^{\frac{2}{3}},t)=\frac{t}{4}-2at^{\frac{2}{3}}+4a^{2}t^{\frac{1}{3}}-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{5}{3}}a)+o(t^{\frac{1}{3}}).$
Then we have
$\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\mathbb{P}\left(2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{5}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-3a^{2}\right.\\ \left.+\min\left\{2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2},0\right\}\geq 0\right).$
Similarly, one can get (5.9) by using (3.15):
$\lim_{t\to\infty}\mathsf{N}_{3}(1,t)\stackrel{{\scriptstyle d}}{{=}}\lim_{t\to\infty}\min\bigl{\{}at^{\frac{2}{3}},2at^{\frac{2}{3}}-\mathcal{N}(2at^{\frac{2}{3}},t)+\mathcal{N}(4at^{\frac{2}{3}},t)\bigr{.}\\ \bigl{.}+\max\{0,at^{\frac{2}{3}}-\mathcal{N}(1,t)+\mathcal{N}(2at^{\frac{2}{3}},t)\}\bigr{\}}.$

### v2

@@
 See Figure 2.
+Let $f(t)$ represent the position of the unique second class particle in the process $\eta_{t}$.
 The second half-line open TASEP (we denote by $\tilde{\eta}_{t}$) is similar with the first one, but with a different initial condition.
@@
 Obviously, the initial conditions $\eta_{0}$ and $\tilde{\eta}_{0}$ create one shock.
+Let $\mathbf{N}_{c}(x,t)$ denote the number of particles of color $c$ in the process $\tilde{\eta}_{t}$ at time $t$ that are located weakly to the right of $x\in\mathbb{R}_{\geq 0}$, where $c=1,2$.
 […]
@@
 The dynamics of $\xi_{t}$ are also given by Figure 2.
+We denote the position of the unique second class particle in the process $\xi_{t}$ by $g(t)$.
 The two-shock case of the second initial configuration $\tilde{\eta}_{0}$ (we denote by $\tilde{\xi}_{0}$) can be obtained from $\xi_{0}$ by replacing the particles on the interval $[N+1,N+M]$ with the second class particles, the particles on the interval $[N+2M+1,2N+2M+1]$ with the third class particles (we denote by 3).
@@
 The dynamics of $\tilde{\xi}_{t}$ are also given by Figure 2, but with $k\in\{2,3,+\infty\}$, $i<j$ and $i,j\in\{1,2,3,+\infty\}$.
+Let $\mathsf{N}_{i}(x,t)$ be the number of particles of color $i$ in $\tilde{\xi}_{t}$ which are weakly to the right of $x\in\mathbb{R}$, where $i\in\{1,2,3\}$.
 […]
@@
 Definition 1.1.
-For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we define the DEHP-tree of $\eta$ by the following way:
+Let $\alpha\in(0,1)$ and $c=\frac{1}{4}$. For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we define the DEHP-tree of $\eta$ by the following way:
   • start from a node represented by a tuple $(\Psi_{1}(\eta),\Phi_{0}(\eta))$, which we call the root.
@@
 The DEHP-tree of $\eta=(0,0,1,0,0,1,1,0,1,1,0,0)$; the DEHP-tree of $\eta$ start from the node $(5,7)$ and end with the nodes $(k,0)$ for $0\leq k\leq 5$; the nodes $(5,7)$, $(5,6)$, $(4,5)$, $(4,4)$, $(2,3)$, $(0,2)$, $(0,1)$ have only one child, other nodes have two children.
-Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
+Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha\leq\frac{1}{2}$:
 Theorem 1.2 (Theorem 4.10).
+Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
+$\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n}.$
+Let $0\leq m\leq M_{2}$, the asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
+$\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}.$
+Our second main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha>\frac{1}{2}$:
+Theorem 1.3 (Theorem 4.11 modulo Conjecture 4.3).
+Let $\alpha>\frac{1}{2}$ and we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
 The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
-(1.1) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n},$
-when $\alpha\leq\frac{1}{2}$, where we use the binomial coefficient $\binom{p}{k}=\frac{p!}{k!(p-k)!}$; and it is given by
-(1.2) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$, where we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
+$\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
 Let $0\leq m\leq M_{2}$, the asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
-(1.3) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}$
-when $\alpha\leq\frac{1}{2}$, and it is given by
-(1.4) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$.
-The asymptotic distribution of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ have the following relationships:
-(1.5) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=M_{2}+1).$
-The asymptotic distribution of the unique second class particle disappearing from $\eta_{t}$ is given by the following formula:
-(1.6) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\sum_{m=0}^{M_{2}}\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m).$
-Our second main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$:
-Theorem 1.3 (Theorem 4.10).
+$\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
+Our third main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha\leq\frac{1}{2}$:
+Theorem 1.4 (4.12).
+Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
+$\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}.$
+Let $0\leq s\leq N$, the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
+$\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}.$
+Our fourth main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha>\frac{1}{2}$:
+Theorem 1.5 (Theorem 4.13 modulo Conjecture 4.3).
+Let $\alpha>\frac{1}{2}$ and we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
 The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
-(1.7) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}$
-when $\alpha\leq\frac{1}{2}$, and it is given by
-(1.8) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$, where we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
+$\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
 Let $0\leq s\leq N$, the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
-(1.9) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}$
-when $\alpha\leq\frac{1}{2}$, and it is given by
-(1.10) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$.
-The asymptotic distribution of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ have the following relationships:
-(1.11) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=N+1).$
-The asymptotic distribution of the unique second class particle disappearing from $\xi_{t}$ is given by the following formula:
-(1.12) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\sum_{s=0}^{N}\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s).$
+$\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
 Remark 1.6.
 The asymptotic distribution of the height function of second class particles in $\tilde{\xi}_{t}$ is similar with (4.13) and (4.15).
-For the sake of brevity, we only provide the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$.
-The third class particles can be regarded as the second-type second class particles.
+For the sake of brevity, we only provide the asymptotic distribution of the height function of second class particles in $\tilde{\xi}_{t}$.
 1.4. Asymptotic distribution of the second class particles under the KPZ scaling
-In this section, we consider the asymptotic probability distribution of the second class particles under the KPZ scaling $M_{1}=M_{2}=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$.
+In this section, we consider the asymptotic probability distribution of the second class particles under the KPZ scaling $t^{\frac{2}{3}}$.
 Below we will use the half-space ${Airy}_{2}$ process $\mathcal{A}_{\varpi}^{HS}$ which was introduced in [BBCS18a], see section 5 for the definition.
@@
 When the clock with rate $\alpha$ rings, a swap occurs between $\eta(-1)$ and $\eta(1)$ if $\eta(-1)<\eta(1)$.
-Otherwise, we remain the configuration of system unchanged.
+Otherwise, we leave the system configuration unchanged.
 See Figure 5 for an illustration.
@@
 Note that by definition $W(0)=T_{id}$, which indicates $\omega_{0}=id$.
+Remark 2.1.
+While the constructions in this section describe stochastic processes on a finite space, they can be extended to the infinite setting using the standard argument of Harris ([Har78]).
+Namely, at any fixed time, the processes on an infinite space are, with probability 1, (infinite) collections of finite-space processes that did not interact with each other.
+Consequently, the results for processes on an infinite space follow directly from the corresponding results for their finite-space counterparts.
+Remark 2.2.
+Since the random element of the Hecke algebra and the corresponding random permutation that it induces are equivalent, for convenience, we will use the two concepts equivalently in the following descriptions and do not mention it again.
 The reason we describe the evolution of colored half-space TASEP as the random walk on the type BC Hecke algebra is that we wish to use a key property of the Hecke algebra $\mathcal{H}(B_{n})$: the anti-involution $\iota$, which sends $\iota(T_{w})=T_{w^{-1}}$ for all $w\in B_{n}$.
@@
 (2.2) $\iota(T_{\omega_{n}}T_{\omega_{n-1}}\cdots T_{\omega_{2}}T_{\omega_{1}})=\iota(T_{\omega_{1}})\iota(T_{\omega_{2}})\cdots\iota(T_{\omega_{n-1}})\iota(T_{\omega_{n}}),$
-which can be straightforwardly proved by induciton in $l(\omega)$ with the use of (2.1).
+which can be straightforwardly proved by induction in $l(\omega)$ with the use of (2.1).
 2.2 was first introduced in probabilistic setting by [Buf20].
@@
 We only prove the first part of above theorem, the proof of the second part is similar. By definition, we can write down the formula of $\eta_{t}^{gen;1}$:
-$\eta_{t}^{gen;1}=W(t)T_{s_{k_{n}}}\cdots T_{s_{k_{1}}}T_{id},$
+$\eta_{t}^{gen;1}=W(t)T_{s_{k_{n}}}\cdots T_{s_{k_{1}}},$
 which is equal to the following formula:
-$\iota(\eta_{t}^{gen;2})=\iota(T_{s_{k_{1}}}\cdots T_{s_{k_{n}}}W(t)T_{id}),$
+$\iota(\eta_{t}^{gen;2})=\iota(T_{s_{k_{1}}}\cdots T_{s_{k_{n}}}W(t)),$
 where we use Lemma 2.3 and Lemma 2.5. ∎
@@
 Let $M_{1},M_{2}$ be positive integers, and consider $\pi_{M_{1},M_{2}}$ be a signed permutation of the set $\{-M_{1}-M_{2}-1,\cdots,-1,1,\cdots,M_{1}+M_{2}+1\}$ such that $\pi_{M_{1},M_{2}}(i)=M_{1}+M_{2}+2-i$ and $\pi(-i)=-\pi(i)$ for $i=1,2,\cdots,M_{1}+M_{2}+1$.
-We assume the length $l(\pi_{M_{1},M_{2}})=m$ and let $\pi_{M_{1},M_{2}}=s_{k_{m}}s_{k_{m-1}}\cdots s_{k_{1}}$ be a minimal length dedomposition of $\pi_{M_{1},M_{2}}$ into the Coxeter generators (there can be many such decompositions, we only choose one of them).
+We assume the length $l(\pi_{M_{1},M_{2}})=m$ and let $\pi_{M_{1},M_{2}}=s_{k_{m}}s_{k_{m-1}}\cdots s_{k_{1}}$ be a minimal length decomposition of $\pi_{M_{1},M_{2}}$ into the Coxeter generators (there can be many such decompositions, we only choose one of them).
 Consider the first process $\eta_{t}^{gen;1}$ introduced in Section 2.3 involving data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$.
@@
 Remark 3.1.
+In this paper, we consider several projections of fully colored half-space TASEP where we map colors $(-\infty,K_{1}]$ to 1, colors $[K_{1}+1,K_{2}]$ to 2, $\ldots$, colors $[K_{N-1}+1,K_{N}]$ to $N$, and colors $[K_{N},+\infty]$ to +$\infty$+ ∞, for some fixed $N\in\mathbb{N}$ and $-\infty<K_{1}<K_{2}<\cdots<K_{N}<+\infty$.
+Note that the set of colors is linearly ordered, and particle with a smaller color interacts with a particle with a larger color as a particle with hole.
+It follows immediately from verifying the marginal transition rates that these projections have the same laws as the multi-species extensions of the half-line open TASEP.
+Remark 3.2.
 Because the dynamics on $\mathbb{N}$ and $-\mathbb{N}$ are symmetric, and the initial configurations (3.1) and (3.2) are filled with first class particles, therefore we can identity half-space TASEP $\eta_{t}$ (resp.
@@
 That’s why we don’t change the notations.
-The reason is the same for the two-shock case in Section 3.4, we will not state it again later.
-Both of above tricolor half-space TASEPs illustrate the second class particles in one shock.
-We aim to investigate the characteristics of the sole second-class particle in the half-space by examining the first process $\eta_{t}$.
-Simultaneously, we will explore the behavior of the multi-colored height functions in the half-space through the analysis of the second process $\tilde{\eta}_{t}$.
-Let $f(t)$ represent the position of the unique second class particle in the process $\eta_{t}$.
-Additionally, let $\mathbf{N}_{c}(x,t)$ denote the number of particles of color $c$ in the process $\tilde{\eta}_{t}$ at time $t$ that are located weakly to the right of $x\in\mathbb{R}_{\geq 0}$, where $c=1,2$, namely,
-(3.3) $\mathbf{N}_{c}(x,t)=\sum_{z\geq x}\delta_{\tilde{\eta}_{t}(z),c},\quad\text{for}\ c=1,2.$
-These quantities will be studied by establishing connections to a more straightforward process using the color-position symmetry.
+The reason is the same for the two-shock case in Section 3.4.
 3.2. Standard half-line open TASEP with step initial condition
@@
 Using the color-position symmetry theorem 2.7, we have
-$\mathbb{P}((\iota\eta)_{t}^{gen;1}(M_{2}+1)<0)=\mathbb{P}(\eta_{t}^{gen;2}(M_{2}+1)<0).$
-Considering the definition of $\eta_{t}^{gen;2}$ and the projection introduced in Section 3.2, the event that the color $\eta_{t}^{gen;2}(M_{2}+1)$ in the position $M_{2}+1$ of $\eta_{t}^{gen;2}$ is less than 0 is equivalent to the event in half-line TASEP (3.4) that the number of particles in the set $\{1,\cdots,M_{1}+M_{2}+1\}$ is at least $M_{1}+1$.
+$\mathbb{P}((\iota\eta)_{t}^{gen;1}(M_{2}+1)<0)=\mathbb{P}(\eta_{t}^{gen;2}(M_{2}+1)<0),$
+where $\eta_{t}^{gen;2}(M_{2}+1)<0$ means the color in the position $M_{2}+1$ is less than 00 in the random configuration $\eta_{t}^{gen;2}$.
+Recall that we get the random configuration $\eta_{t}^{gen;2}$ by the following way: we start from the identity configuration, and then we perform the continuous time process for time $t$, finally we sort the colors on the interval $[1,M_{1}+M_{2}+1]$ in a decreasing order, at the same time we sort the colors on the interval $[-M_{1}-M_{2}-1,-1]$ in a decreasing order.
+Consider the projection introduced in Section 3.2: we say that if a color is $\leq-1$, then we have a particle; while if a color is $\geq 1$, then we have a hole.
+Under this projection, we can interpret $\eta_{t}^{gen;2}$ as the standard half-line open TASEP: at time $t=0$, we have the step initial condition; after running the dynamics for time $t$, we rearrange the holes on the interval $[1,M_{1}+M_{2}+1]$ to the left and the particles to the right.
+Thus, the event $\eta_{t}^{gen;2}(M_{2}+1)<0$ is equivalent to the event in half-line TASEP (3.3) that the number of particles in the set $\{1,\cdots,M_{1}+M_{2}+1\}$ is at least $M_{1}+1$.
 Namely, we have the equation 3.5.
@@
 Proof.
-We use Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$ introduced in Section 3.1 in the folloing. By the relation of $\tilde{\eta}_{t}$ and $\eta_{t}^{gen;1}$, we have
+We use Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$ introduced in Section 3.1 in the following. By the relation of $\tilde{\eta}_{t}$ and $\eta_{t}^{gen;1}$, we have
 $\left(\mathbf{N}_{1}(1,t),\mathbf{N}_{2}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\#\{i:i<0,\iota(\eta_{t}^{gen;1})(i)\geq 1\},\#\{i:1\leq i\leq M_{2}+1,\iota(\eta_{t}^{gen;1})(i)\geq 1\}\right),$
@@
 (#$\left(\#\{i:i<0,\iota(\eta_{t}^{gen;1})(i)\geq 1\},\#\{i:1\leq i\leq M_{2}+1,\iota(\eta_{t}^{gen;1})(i)\geq 1\}\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\},\#\{i:1\leq i\leq M_{2}+1,\eta_{t}^{gen;2}(i)\geq 1\}\right),$
-Similar with the proof of Theorem 3.2, we can get 3.11. ∎
+Similar with the proof of Theorem 3.3, we can get 3.7. ∎
 Remark 3.5.
@@
 We can generalize the results in Theorem 3.3 and Theorem 3.4 to the distribution identities of distribution functions by introducing the half-line open TASEP with the shifted step initial condition(denoted by $\eta_{t}^{step(y)}$, $y\in\mathbb{N}\cup-\mathbb{N}$) :
-(3.9) $\eta_{0}^{step(y)}(x)$ =$=\begin{cases}1,&\text{ if }x\leq y,\\ +\infty,&\text{ if }x>y.\end{cases}$
+$\eta_{0}^{step(y)}(x)$ =$=\begin{cases}1,&\text{ if }x\leq y,\\ +\infty,&\text{ if }x>y.\end{cases}$
 where $x\in\mathbb{N}\cup-\mathbb{N}$.
@@
 For any $x\in\mathbb{R}$, we use $\mathcal{N}^{y}(x,t)$ to represent the number of particles which are weakly to the right of $x$ in $\eta_{t}^{step(y)}$, namely,
-(3.10) $\mathcal{N}^{y}(x,t)=\sum_{z\geq x}\delta_{\eta_{t}^{step(y)}(z),1}.$
+$\mathcal{N}^{y}(x,t)=\sum_{z\geq x}\delta_{\eta_{t}^{step(y)}(z),1}.$
 Then the equation (3.5) can be generalized to the following form:
@@
 And the equation (3.7) can be generalized to the following form:
-(3.11) $\left(\mathbf{N}_{1}(y,t),\mathbf{N}_{2}(y,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\mathcal{N}^{y}(1,t)-y,\min\left\{M_{1}+M_{2}+1-\mathcal{N}^{y}(1,t)+\mathcal{N}^{y}(M_{1}+M_{2}+2,t),M_{2}+1\right\}\right).$
-Since we know very little about the asymptotic properties of $\mathcal{N}^{y}(x,t)$, this method is far to reach.
+$\left(\mathbf{N}_{1}(y,t),\mathbf{N}_{2}(y,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\mathcal{N}^{y}(1,t)-y,\min\left\{M_{1}+M_{2}+1-\mathcal{N}^{y}(1,t)+\mathcal{N}^{y}(M_{1}+M_{2}+2,t),M_{2}+1\right\}\right).$
+We will address the asymptotic analysis in our future work.
 Remark 3.7.
@@
 (3.9) $\tilde{\xi}_{0}(x)$ =$=\begin{cases}1,&\text{ if }x\leq-1,\\ 2,&\text{ if }N+1\leq x\leq N+M,\\ 3,&\text{ if }2M+N+1\leq x\leq 2M+2N+1,\\ +\infty,&\text{otherwise}.\end{cases}$
-We denote the position of the unique second class particle in the process $\xi_{t}$ by $g(t)$.
-Further, let $\mathsf{N}_{i}(x,t)$ be the number of particles of color $i$ in $\tilde{\xi}_{t}$ which are weakly to the right of $x\in\mathbb{R}$, where $i\in\{1,2,3\}$.
+For the same reason as in Remark 3.2, the model $\xi_{t}$ is equivalent to the half-line open TASEP shown in the upper part of Figure 3, and the model $\tilde{\xi}_{t}$ is equivalent to the one shown in the bottom part of Figure 3.
+We use the same notations $g(t)$ and $\mathsf{N}_{i}(x,t)$ as defined in Section 1.2.
 Theorem 3.8.
@@
 Proof.
-Let $\pi_{1}$ be a signed permutation of the set $\{-2M-2N-1,\cdots,-M-1,M+1,\cdots,2M+2N+1\}$ such that $\pi_{1}(M+1+i)=\pi_{1}(2M+2N+1-i)$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+2N$.
-And let $\pi_{2}$ be a signed permutation of the set $\{-M-N,\cdots,-1,1,\cdots,M+N\}$ such that $\pi_{2}(1+i)=\pi_{1}(M+N-i)$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+N-1$.
+Let $\pi_{1}$ be a signed permutation of the set $\{-2M-2N-1,\cdots,-M-1,M+1,\cdots,2M+2N+1\}$ such that $\pi_{1}(M+1+i)=2M+2N+1-i$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+2N$.
+And let $\pi_{2}$ be a signed permutation of the set $\{-M-N,\cdots,-1,1,\cdots,M+N\}$ such that $\pi_{2}(1+i)=M+N-i$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+N-1$.
 Consider a minimal length decomposition of permutation $\pi_{2}\cdot\pi_{1}$ into Coxeter generators: $\pi_{2}\cdot\pi_{1}=s_{\ell_{p}}s_{\ell_{p-1}}\cdots s_{\ell_{1}}$, where $l(\pi_{2}\cdot\pi_{1})=p$.
@@
 We can get $\xi_{t}$ from $\eta_{t}^{gen;1}$(with Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$) by identifying the particle with color $M+N+1$ as the second class particle (denote by 2), particles with colors less than $M+N+1$ as first class particles (denote by 1), and particles with colors greater than $M+N+1$ are holes (denote by +$\infty$+ ∞).
-That means:
+This implies:
 $\mathbb{P}(g(t)<0)=\mathbb{P}(\iota(\eta_{t}^{gen;1})(M+N+1)<0).$
@@
 and the following relation:
-$\left(\mathsf{N}_{1}(1,t),\mathsf{N}_{2}(1,t),\mathsf{N}_{3}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\left(\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\},\right.\\ \left.\#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\},\#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\right),$
-where we use Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct $\eta_{t}^{gen;2}$(recall the definition introduced in Section 2.3).
-Considering the definition of $\eta_{t}^{gen;2}$ and the projection introduced in Section 3.2, the color of the particle in position $M+N+1$ is less than 0 in $\eta_{t}^{gen;2}$ is equivalent to the event in half-line TASEP 3.4 that the number of particles in the set $\{M+1,\cdots,2M+2N+1\}$ is at least $M+N+1$.
+(3.12) $\left(\mathsf{N}_{1}(1,t),\mathsf{N}_{2}(1,t),\mathsf{N}_{3}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\left(\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\},\right.\\ \left.\#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\},\#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\right),$
+where we use Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct $\eta_{t}^{gen;2}$.
+Recall that we get the random configuration $\eta_{t}^{gen;2}$ by the following way: we start from the identity configuration; and then we perform the continuous time process for time $t$; after that we separately sort the colors on the interval $[1,M+N]$ and $[-M-N,-1]$ in a decreasing order at the same time; finally we separately sort the colors on the interval $[M+1,2M+2N+1]$ and $[-2M-2N-1,-M-1]$ in a decreasing order.
+Consider the projection introduced in Section 3.2: we say that if a color is $\leq-1$, then we have a particle; while if a color is $\geq 1$, then we have a hole.
+Under this projection, we can interpret $\eta_{t}^{gen;2}$ as the standard half-line open TASEP: at time $t=0$, we have the step initial condition; after running the dynamics for time $t$, we rearrange the holes on the interval $[1,M+N]$ to the left and the particles to the right; finally, we rearrange the holes on the interval $[M+1,2M+2N+1]$ to the left and the particles to the right.
+Thus, the event $\eta_{t}^{gen;2}(M+N+1)<0$ is equivalent to the event in half-line TASEP (3.3) that the number of particles $\mathcal{N}(N+M+1,t)-\mathcal{N}(2N+2M+2,t)+\min\{N,\mathcal{N}(1,t)-\mathcal{N}(N+M+1,t)\}$ are at least $M+N+1$.
 Namely, we have the equation 3.10.
-Similarly, we have the following distribution identity:
-#$\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \#\{i:i>0,\eta_{t}^{gen;2}(i)\leq-1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \mathcal{N}(1,t),\\ \#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \min\{M,M+N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\\ \#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\stackrel{{\scriptstyle d}}{{=}}\min\bigl{\{}N+1,M+N+1-\mathcal{N}(M+N+1,t)\bigr{.}\\ \bigl{.}+\mathcal{N}(2M+2N+2,t)+\max\{0,N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\bigr{\}},$
-which lead to the proof of 3.15. ∎
+Under the above projection, we can also establish the following identities between the events in $\eta_{t}^{gen;2}$ and the events in standard half-line open TASEP:
+#$\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\}=\#\{i:i>0,\eta_{t}^{gen;2}(i)\leq-1\}\ =\mathcal{N}(1,t),\\ \#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\}=\min\{M,M+N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\\ \#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}=\min\bigl{\{}N+1,M+N+1-\mathcal{N}(M+N+1,t)\bigr{.}\\ \bigl{.}+\mathcal{N}(2M+2N+2,t)+\max\{0,N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\bigr{\}},$
+which lead to the proof of 3.11 by using the joint distribution identity (3.12). ∎
 4. Asymptotic distribution of second class particles under constant scaling
@@
 4.1. Stationary measure of half-line open TASEP
-According to Liggett’s results on half-line open ASEP [Ligg75], we can characterize the stationary measure of half-line TASEP with step initial condition (denoted by $\eta_{t}^{step}$) as follows:
+Let the parameter $\alpha\in(0,1)$ represent the particle density imposed by the reservoir at site 1.
+When $\alpha\leq\frac{1}{2}$, Liggett’s result [Ligg75, Theorem 1.8] characterizes the stationary distribution of the half-line open TASEP with step initial condition (denoted by $\eta_{t}^{step}$) as follows:
 Theorem 4.1.
-Let parameter $\alpha\in(0,1)$ corresponds to the density of particles that the reservoir imposes at site 1.
-When $\alpha\leq\frac{1}{2}$, the process $\eta_{t}^{step}$ admits stationary measures that are product i.i.d.
-Bernoulli with parameter $\alpha$.
-When $\alpha>\frac{1}{2}$, stationary measures (denoted by $\mu_{1/2}^{\alpha}$) of process $\eta_{t}^{step}$ are spatially correlated and more complicated: there is a rarefaction fan with a density of particles $\alpha$ near the origin and density $\frac{1}{2}$ at +$\infty$+ ∞.
-Following the seminal work of [DEHP93], Grosskinsky [Gro04] has given a explicit characterisation of the stationary measure $\mu_{1/2}^{\alpha}$ of half-line open TASEP with a matrix product ansatz:
+For $\alpha\leq\frac{1}{2}$, the half-line open TASEP with step initial condition $\eta_{t}^{\mathrm{step}}$ converges to a product measure of i.i.d.
+Bernoulli random variables with parameter $\alpha$.
+When $\alpha>\frac{1}{2}$, Liggett [Ligg75] showed that the limiting stationary measures of the process $\eta_{t}^{step}$ are spatially correlated and have an asymptotic density of $\frac{1}{2}$.
+Based on physical arguments, these limiting stationary measures are conjectured to coincide with the stationary measure described in [Gro04, Theorem 3.2] with stationary current $c=\frac{1}{4}$.
+However, establishing this connection rigorously remains an open and technically challenging problem.
+In this paper, we present several theorems that are conditional on this conjecture.
+For completeness, we state below the stationary measure given in [Gro04, Theorem 3.2] for the case $c=\frac{1}{4}$ and the above conjecture.
 Theorem 4.2.
-The $L$ dimensional distributions of stationary measures $\mu_{1/2}^{\alpha}$ for the process $\eta_{t}^{step}$ are given by the following matrix product ansatz:
-(4.1) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{L}=\eta_{L}\}=\left\langle w\right|\prod_{x=1}^{L}\left(\eta_{x}D+(1-\eta_{x})E\right)\left|v\right\rangle,$
-where $L\in\mathbb{N}$, $(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, the nonnegative matrices $D$, $E$ and vectors $\left\langle w\right|$ and $\left|v\right\rangle$ satisfy the following algebraic relations with $c=\frac{1}{4}$:
+Let $D$, $E$ be non-negative matrices, and $\left\langle w\right|$, $\left|v\right\rangle$ be vectors satisfying the following algebraic relations with $c=\frac{1}{4}$:
 (4.1) $DE=c(D+E),$
@@
 (4.3) $(D+E)\left|v\right\rangle=\left|v\right\rangle,$
-(4.5) $\left\langle w\right|\left|v\right\rangle=1.$
+(4.4) $\langle w|v\rangle=1.$
+Then for $\alpha>\frac{1}{2}$, the probability measure $\mu_{1/2}^{\alpha}$ defined via the following matrix product ansatz:
+(4.5) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{L}=\eta_{L}\}=\left\langle w\right|\prod_{x=1}^{L}\left(\eta_{x}D+(1-\eta_{x})E\right)\left|v\right\rangle,$
+for any $L\in\mathbb{N}$ and $(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, is stationary for the dynamics of half-line open TASEP.
+Conjecture 4.3.
+For $\alpha>\frac{1}{2}$, the half-line open TASEP with step initial condition $\eta_{t}^{\mathrm{step}}$ converges to the probability measure $\mu_{1/2}^{\alpha}$ defined in 4.5.
+The main result of this section is the following Theorem.
+Theorem 4.4.
+For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we define the DEHP-tree of $\eta$ by definition 1.1.
+The stationary measure 4.5 has the explicit combinatorial formula:
+(4.6) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),$
+where $\Psi_{1}(\eta)$ and $Z_{k}(\eta)$ are defined in Section 1.3.
+In the rest of this section, we will analyze the structure of matrix product ansatz (4.5) and give a proof of theorem 4.4.
 With rules (4.1), (4.2) and (4.4), the probability of any cylinder set $(\eta_{1},\cdots,\eta_{L})$ can be directly written as a linear combination of $\left\langle w\right|D^{k}\left|v\right\rangle$:
@@
 $=$ $\left\langle w\right|\underbrace{E\ldots E}_{\tau_{0}}\underbrace{D\ldots D}_{\sigma_{1}}\underbrace{E\ldots E}_{\tau_{1}}\underbrace{D\ldots D}_{\sigma_{2}}\cdots\underbrace{E\ldots E}_{\tau_{n-1}}\underbrace{D\ldots D}_{\sigma_{n}}\underbrace{E\ldots E}_{\tau_{n}}\left|v\right\rangle$
-$=$ $(c/\alpha)^{\tau_{0}}\left\langle w\right|\underbrace{D\ldots D}_{\sigma_{1}}\underbrace{E\ldots E}_{\tau_{1}}\underbrace{D\ldots D}_{\sigma_{2}}\cdots\underbrace{E\ldots E}_{\tau_{n-1}}\underbrace{D\ldots D}_{\sigma_{n}}\underbrace{E\ldots E}_{\tau_{n}}\left|v\right\rangle$
-::=:= $(c/\alpha)^{\tau_{0}}F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n}),$
-where we use the algebraic relation (4.3) and the expression (4.7) of $\eta$.
-Our main task is to express $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ as a linear combination of $\left\langle w\right|D^{k}\left|v\right\rangle$.
+::=:= $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n}),$
+Our main task is to express $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ as a linear combination of $\left\langle w\right|D^{k}\left|v\right\rangle$.
 As an appetizer, we first give the explicit combinatorial formula of $F_{1}(\sigma_{1},\tau_{1})$ with only one pair of blocks $\underbrace{D\ldots D}_{\sigma_{1}}\underbrace{E\ldots E}_{\tau_{1}}$.
@@
 $(x,y)\iff\underbrace{D\ldots D}_{x}\underbrace{E\ldots E}_{y}$
-Now we are in a position to extend the definition of one-layer DEHP-tree to $n$-layer DEHP-tree, and then give the explicit combinatorial formula of $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ with arbitrary $n$ pairs of blocks $\prod_{i=1}^{n}\underbrace{D\ldots D}_{\sigma_{i}}\underbrace{E\ldots E}_{\tau_{i}}$.
-We introduce the notation $\tilde{\Psi}_{i}=\sum_{j=1}^{i}\sigma_{j}$ and $\tilde{\Phi}_{i}=\sum_{j=1}^{i}\tau_{j}$.
+Now we are in a position to extend the definition of one-layer DEHP-tree to multi-layer DEHP-tree, and then give the explicit combinatorial formula of $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ with arbitrary pairs of blocks $\prod_{i=0}^{n}\underbrace{D\ldots D}_{\sigma_{i}}\underbrace{E\ldots E}_{\tau_{i}}$, where $\sigma_{0}=0$.
+We introduce the notation $\tilde{\Psi}_{i}=\sum_{j=0}^{i}\sigma_{j}$ and $\tilde{\Phi}_{i}=\sum_{j=0}^{i}\tau_{j}$.
 Definition 4.7.
-The $n$-layer DEHP-tree of $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$ is a rooted tree with DEHP weight, which is defined by the following recursive way:
-  • start from an one-layer DEHP-tree of $(\tilde{\Psi}_{1},\tilde{\Phi}_{1})$;
+The multi-layer DEHP-tree of $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$ is a rooted tree with DEHP weight, which is defined by the following recursive way:
+  • start from an one-layer DEHP-tree of $(\tilde{\Psi}_{0},\tilde{\Phi}_{0})$;
   • construct the $i$-layer DEHP-tree of $(\tilde{\Psi}_{i},\tilde{\Phi}_{i})$ from the $(i-1)$-layer DEHP-tree of $(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})$: we add $(\sigma_{i},\tau_{i})$ to every node of the $(i-1)$-layer DEHP-tree of $(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})$ and then continue the tree construction by performing the generation of one-layer DEHP-tree of $(\tilde{\Psi}_{i},\tau_{i})$.
@@
 Proposition 4.8.
-We can give the explicit formula of $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ by the above $n$-layer DEHP-tree:
-(4.9) $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})=\sum_{k=0}^{\tilde{\Psi}_{n}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})\left\langle w\right|D^{k}\left|v\right\rangle.$
+We can give the explicit formula of $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ by the above multi-layer DEHP-tree:
+(4.10) $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})=\sum_{k=0}^{\tilde{\Psi}_{n}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})\left\langle w\right|D^{k}\left|v\right\rangle.$
 Proof.
-We can conclude the proof from the definition 4.5. ∎
+By definition, we have $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})=\left\langle w\right|\prod_{j=0}^{n}\underbrace{D\ldots D}_{\sigma_{j}}\underbrace{E\ldots E}_{\tau_{j}}\left|v\right\rangle$, where $\sigma_{0}=0$.
+We calculate the product step by step.
+First, $\left\langle w\right|\underbrace{E\ldots E}_{\tau_{0}}$ is given by the one-layer DEHP-tree: $Z_{0}(0,\tau_{0})\left\langle w\right|$.
+If we already have $\left\langle w\right|\prod_{j=0}^{i-1}\underbrace{D\ldots D}_{\sigma_{j}}\underbrace{E\ldots E}_{\tau_{j}}=\sum_{k=0}^{\tilde{\Psi}_{i-1}}Z_{k}(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})\left\langle w\right|D^{k}$, then $\left\langle w\right|\prod_{j=1}^{i}\underbrace{D\ldots D}_{\sigma_{j}}\underbrace{E\ldots E}_{\tau_{j}}$ can be expressed by
+$\sum_{k=0}^{\tilde{\Psi}_{i-1}}Z_{k}(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})\left\langle w\right|D^{k}\underbrace{D\ldots D}_{\sigma_{i}}\underbrace{E\ldots E}_{\tau_{i}}.$
+By the proof of Proposition 4.6 and the definition of $i$-layer DEHP-tree, the above expression is equivalent to $\sum_{k=0}^{\tilde{\Psi}_{i}}Z_{k}(\tilde{\Psi}_{i},\tilde{\Phi}_{i})\left\langle w\right|D^{k}$.
+Finally, at step $n$, we can get the result (4.10).
+∎
 $(3,3)$$c$$(3,2)$$c$$(2,3)$$c$$(3,1)$$c$$(2,2)$$c$$(3,0)$$c$$(2,1)$$(2,0)$$c$$c$$(1,1)$$c$$c$$(1,0)$$(0,1)$$(0,0)$$c/\alpha$$c/\alpha$$c/\alpha$$(5,5)$$c$$(5,4)$$c$$(4,5)$$c$$(5,3)$$c$$(4,4)$$c$$(5,2)$$c$$(4,3)$$c$$(5,1)$$c$$(4,2)$$c$$c$$(3,3)$$c$$c$$c$$c$$c$$c$$(5,0)$$(4,1)$$(3,2)$$(2,3)$$c$$c$$c$$c$$(4,0)$$(3,1)$$(2,2)$$c$$c$$c$$c$$(3,0)$$(2,1)$$(1,2)$$c$$c$$c$$c$$(2,0)$$(1,1)$$(0,2)$$c$$c$$(1,0)$$(0,1)$$c/\alpha$$c/\alpha$$c/\alpha$$c/\alpha$$c/\alpha$$(0,0)$
@@
 Corollary 4.9.
-For any $\eta$ with the form 4.7, the stationary measure 4.1 has the explicit formula:
-(4.10) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\tilde{\Psi}_{n}}\frac{c^{\tau_{0}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})}{2^{k}\alpha^{\tau_{0}}}\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),$
-where $c=\frac{1}{4}$ and $Z_{k}(\Psi_{n},\Phi_{n})$ is given by the partition function of $(k,0)$ in the $n$-layer DEHP-tree.
-Remark 4.8.
-One can check the construction of DEHP-tree in Definition 1.1 is equivalent to the above construction.
-The difference is that we incorporate the term $\left\langle w\right|\underbrace{E\ldots E}_{\tau_{0}}$ into the Definition 1.1.
-The reason we do this is that the expression of formula is more neat:
-(4.11) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
-where $\Psi_{1}(\eta)$ and $Z_{k}(\eta)$ are defined in Section 1.3.
+For any $\eta$ with the form 4.8, the stationary measure 4.5 has the explicit formula:
+(4.11) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\tilde{\Psi}_{n}}\frac{Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})}{2^{k}}\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),$
+where $Z_{k}(\Psi_{n},\Phi_{n})$ is given by the partition function of $(k,0)$ in the multi-layer DEHP-tree of $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$.
+Now we can provide the proof of theorem 4.4.
+Proof.
+In the above, we use the recursive way to give the explicit formula of stationary measure 4.5.
+Now we use the notation in Section 1.3 to give a direct way.
+In definition 4.7, we construct the multi-layer DEHP-tree by starting from node $(\tilde{\Psi}_{0},\tilde{\Phi}_{0})$, which is exactly the starting node $(\Psi_{1}(\eta),\Phi_{0}(\eta))$ in definition 1.1.
+By the definition 4.5, any node in the multi-layer DEHP-tree 4.7 have either one child or two children.
+When $(x,y)\in\{(x,y):x=\Psi_{i}(\eta)\ \text{and}\ \Phi_{i}(\eta)<y\leq\Phi_{i-1}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)+1\}$, the node $(x,y)$ has only one child.
+When $(x,y)\in\{(x,y):x=\Psi_{i}(\eta)\ \text{and}\ 1\leq y\leq\Phi_{i}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)\}\cup\{(x,y):\Psi_{i+1}(\eta)<x<\Psi_{i}(\eta),1\leq y\leq\Phi_{i}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)\}$, the node $(x,y)$ has two children.
+One can easily find that the above observations are exactly the rules in definition 1.1.
+The proof is done.
+∎
 4.2. Explicit formulas for asymptotic distribution of the second class particles
-Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
+Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha\leq\frac{1}{2}$:
 Theorem 4.10.
+Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
+(4.12) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n}.$
+Let $0\leq m\leq M_{2}$, the asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
+(4.13) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}.$
+Proof.
+The proof directly follows from Theorem 3.3, Theorem 3.4, Theorem 4.1. ∎
+Our second main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha>\frac{1}{2}$:
+Theorem 4.11 (Modulo Conjecture 4.3).
+Let $\alpha>\frac{1}{2}$ and we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
 The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
-(4.12) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n}$
-when $\alpha\leq\frac{1}{2}$, and it is given by
-(4.13) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$, where we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
+(4.14) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
 Let $0\leq m\leq M_{2}$, the asymptotic distribution of the height function of second class particles in $\tilde{\eta}_{t}$ is given by:
-(4.14) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}$
-when $\alpha\leq\frac{1}{2}$, and it is given by
-(4.15) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$.
-The asymptotic distribution of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ have the following relationships:
-(4.16) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=M_{2}+1).$
-The asymptotic distribution of the unique second class particle disappearing from $\eta_{t}$ is given by the following formula:
-(4.17) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\sum_{m=0}^{M_{2}}\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m).$
+(4.15) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
 Proof.
-The proof directly follows from Theorem 3.2, Theorem 3.3, Theorem 4.2 and Remark 4.8. ∎
-Our second main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$:
+The proof directly follows from Theorem 3.3, Theorem 3.4, Theorem 4.2, Conjecture 4.3 and Theorem 4.4. ∎
+Our third main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha\leq\frac{1}{2}$:
 Theorem 4.12.
+Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
+(4.16) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}.$
+Let $0\leq s\leq N$, the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
+(4.17) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}.$
+Proof.
+The proof directly follows from Theorem 3.8, theorem 4.1. ∎
+Our fourth main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha>\frac{1}{2}$:
+Theorem 4.13 (Module Conjecture 4.3).
+Let $\alpha>\frac{1}{2}$ and we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
 The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
-(4.18) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}$
-when $\alpha\leq\frac{1}{2}$, and it is given by
-(4.19) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$, where we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
+(4.18) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
 Let $0\leq s\leq N$, the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$ is given by:
-(4.20) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}$
-when $\alpha\leq\frac{1}{2}$, and it is given by
-(4.21) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
-when $\alpha>\frac{1}{2}$.
-The asymptotic distribution of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ have the following relationships:
-(4.22) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=N+1).$
-The asymptotic distribution of the unique second class particle disappearing from $\xi_{t}$ is given by the following formula:
-(4.23) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\sum_{s=0}^{N}\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s).$
+(4.19) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
 Proof.
-The proof directly follows from Theorem 3.7, Theorem 4.2 and Remark 4.8. ∎
+The proof directly follows from Theorem 3.8, Theorem 4.2, Conjecture 4.3 and Theorem 4.4. ∎
 5. Asymptotic distribution of second class particles under KPZ-type scaling
@@
 (5.1) $\mathrm{Pf}\big{(}J+K\big{)}_{\mathbb{L}^{2}(\mathbb{X},\mu)}=1+\sum_{k=1}^{\infty}\frac{1}{k!}\int_{\mathbb{X}}\dots\int_{\mathbb{X}}\mathrm{Pf}\Big{(}K(x_{i},x_{j})\Big{)}_{i,j=1}^{k}\mathrm{d}\mu^{\otimes k}(x_{1}\dots x_{k}),$
-provided the series converges, and we recall that for an skew-symmetric $2k\times 2k$ matrix $A$, its Pfaffian is defined by
+provided the series converges, and we recall that for a skew-symmetric $2k\times 2k$ matrix $A$, its Pfaffian is defined by
 (5.2) $\mathrm{Pf}(A)=\frac{1}{2^{k}k!}\sum_{\sigma\in\mathcal{S}_{2k}}\mathrm{sign}(\sigma)a_{\sigma(1)\sigma(2)}a_{\sigma(3)\sigma(4)}\dots a_{\sigma(2k-1)\sigma(2k)}.$
@@
 where $o(t^{\frac{1}{3}})$ means that after dividing by $t^{\frac{1}{3}}$ this term will converge to 0 in probability. Therefore, in the $t\to\infty$ limit 5.7 converges to
-$\mathbb{P}\left(2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}t^{\frac{1}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2}\geq 0\right)$
+$\mathbb{P}\left(2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(2^{\frac{2}{3}}a)-2^{-\frac{4}{3}}\mathcal{A}_{\varpi}^{HS}(0)-a^{2}\geq 0\right)$
 then we have (5.5). Similarly, one can get (5.6) by using Theorem 3.4:
@@
 Theorem 5.5.
-Assume that $M=N=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$.
+Assume that $M_{1}=M_{2}=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$.
 Setting $\alpha=\frac{1+2^{4/3}\varpi t^{-1/3}}{2}$, where $\varpi\in\mathbb{R}$.